### v1

# Picocavity-enhanced Raman Spectroscopy of Physisorbed H2 and D2 Molecules

Abstract
We report on tip-enhanced Raman scattering (TERS) of H2 and D2 molecules physisorbed within a plasmonic picocavity at a cryogenic temperature (10 K). The intense Raman peaks resulting from the rotational and vibrational transitions are observed at sub-nanometer gap distances of the junction formed by a Ag tip and Ag(111) surface. We clarify that the predominant contribution of the electromagnetic field enhancement of the picocavity to the detection of a single hydrogen molecule. The gap-distance dependent TERS reveals not only the evolution of the picocavity field, but also the interaction between the molecule and tip/surface, which exhibit nontrivial isotope effects. A significant red-shift and peak broadening of the H–H stretching as the gap distance decreases, while the D–D stretching mode is unaffected. A combination of density functional theory and reduced-dimension models reveals that a distinct anharmonicity in the mode potential of H2 is one cause of the anomalous red-shift, whereas D2 has less anharmonicity due to the geometric isotope effect.
The adsorption of hydrogen molecules on solid surfaces is the first step in permeation for fuel storage [1], catalytic hydrogenation [2], hydrogen embrittlement [3], and nuclear-spin isomer conversion [4]. To understand and control the elementary processes in such physical and chemical phenomena, hydrogen adsorption has been intensively studied in surface science [5, 6, 7]. However, only a few methods can be used to characterize physisorption systems, such as hydrogen on coinage-metal surfaces [8, 9, 10, 11, 12, 13], because cryogenic-temperature measurements are required due to the faint adsorption energy ( meV). Low-temperature scanning tunneling microscopy (LT-STM) has been employed to investigate the dynamics of hydrogen molecules weakly adsorbed on surfaces [14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. Several studies have assigned characteristic features in conductance spectra to the rotational mode of hydrogen in the tip–sample junction [17, 18, 19, 20]. However, the interpretation of the conductance spectra remains controversial; the features within the corresponding energy range [30–60 meV (240–480 cm-1) for H2] could be attributed to configuration switches [14, 24, 25], to a molecule?substrate stretching mode [26], or to multiple phonon excitation [27]. Additionally, LT-STM studies have not provided information on the interatomic stretching mode of the molecule. Therefore, an alternative approach is required to identify the rotational/vibrational state.
Raman spectroscopy is a promising technique for characterizing hydrogen molecules on surfaces since the rotational and vibrational transitions of a homonuclear diatomic molecules are observable. However, the detection demands exceptional sensitivity due to the intrinsically small cross-section. Recent advancements in tip-enhanced and surface-enhanced Raman spectroscopy (TERS/SERS) have enabled the detection of single molecules [28, 29, 30, 31]. Picocavities, atomic-scale protrusions present in plasmonic nanojunctions, play a crucial role in achieving extreme confinement of the electromagnetic field to Ångström scales [32, 33, 31], resulting in a significant enhancement of Raman scattering and ultrahigh spatial resolution [34, 35, 36, 37]. So far, the exceptional sensitivity in TERS/SERS has been demonstrated primarily for organic molecules chemisorbed onto the substrate, where the enhancement benefits from chemical effects arising from the molecule–surface (and molecule–tip) interactions [38]. Therefore, applying TERS/SERS to physisorbed systems is not straightforward, as it depends on the fundamental enhancement mechanisms of Raman scattering in plasmonic picocavities.
Here we demonstrate that LT-TERS can observe single hydrogen molecules physisorbed onto an Ag(111) surface with an Ag tip [Fig. 1(a), inset], identifying their rotational and vibrational states. Furthermore, precise gap-distance control enables us to investigate not only how the picocavity field contributes to the Raman scattering process but also how the subtle van der Waals (vdW) interactions between the molecule and the tip/surface affect the anharmonicity of the potential energy surfaces and the accompanying nuclear quantum effects.
The LT-TERS experiments were performed in an ultra-high vacuum chamber, as described in the Supplemental Material (SM) [39]. The 532-nm incident laser beam was linearly polarized along the tip axis and focused on the apex of a focused-ion-beam-sharpened Ag tip [40] using an in-situ parabolic mirror, and the scattering light was detected by an ex-situ spectrometer. We performed density functional theory (DFT) calculations [41, 42] on a model geometry setup. We considered H2/D2 on a Ag(111) (periodic) surface, employing the Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional and screened pairwise vdW interactions [43] (see SM [39] for details). The preferred adsorption sites (fcc/hcp hollow) and binding energies (30 meV) were in very good agreement with previous studies of isolated molecules on the surface [44, 45]. A single H2/D2 molecule proved not to be stable in our simulations, due to the extremely low diffusion barrier. Therefore, we considered a coverage of 0.69 monolayer (ML), motivated by the results of path-integral molecular dynamics simulations at LT, where a 1 ML coverage was also unstable (see Section I-B in SM [39]). In 0.69 ML, the intermolecular attractive interaction increases the binding energy by 55 meV (Fig. S3 [39]). Over the sample, we put a Ag19 cluster in a cleaved pyramidal geometry as a Ag tip (see Section I-B in SM [39] and Refs. [46, 36] for the tip structure details).
H2 or D2 gas was dosed into the chamber where a clean Ag(111) surface was cooled at 10 K. Regardless of laser illumination, hydrogen molecules on the surface are invisible with LT-STM (Fig. S1 [39]). In contrast, clear peaks were observed in TERS spectra under these conditions. Figure 1(a) displays a typical TERS spectrum of H2 (D2) on Ag(111), showing two intense peaks at 351 (169) and 4121 (2967) cm-1 together with a weak peak at 4443 (3133) cm-1. The three peaks in the TERS spectra can be assigned to rotational/vibrotational modes of hydrogen. Note that broad or asymmetric peak shapes are artifacts in the wide wavenumber range spectra (see Section I-A in SM [39]); higher resolution spectra will be shown later. The structure at 100 cm-1 common to the two spectra originates from Ag phonon modes of the tip [47, 36]. Figure 1(b) depicts the energy diagram of rotational/vibrational levels and possible transitions under the Raman selection rule [the change of the vibrational (rotational) quantum numbers () from the initial to final states should be 0 or (0 or )]. The observed peaks, in order from lowest to highest, correspond to the lowest-energy rotational transition [; “rot” labeled in Fig. 1(a)], the lowest-energy vibrational transition [; “vib”], and their combination [; “rot+vib”], respectively. The three modes were also detected in high-resolution electron energy loss spectroscopy (HREELS) of hydrogen on a polycrystalline Ag film [8] and a Cu(001) surface [9] at approximately 10–12 K; however, better energy resolution (10 cm-1) in TERS than that of the HREELS (28 cm-1 [9]) makes the red-shifts from gas-phase hydrogen more discernible (Table S2). We should also note that only the stable nuclear spin isomer of each isotope, i.e., para-H2 and ortho-D2 [ = even; Fig. 1(b)], is detected in TERS (Fig. S2). On the surface, the stable isomer is dominant due to the ortho–para conversion [8, 4, 44] (see SM [39] for details).
We calculate the H–H bonding distance to be 75.4 pm on Ag(111) (cf. 75.1 pm in the gas phase). Considering a rigid-rotor approximation, the rotational transition energy of para-H2 is calculated to be of 353 cm-1 on the surface, compared to 356 cm-1 in the gas phase. This is in good agreement with the experimental TERS peak at 351 cm-1 (Fig. 1a) slightly red-shifted from 354 cm-1 in the gas phase [48]. Notably, the H–H bonding distance of H2 on Ag(111) varied by only 0.06 pm at the closest tip position with respect to that without any tip included in our simulations. This suggests that the hydrogen molecules on the surface and in the STM junction behave as nearly free rotors.
In the harmonic approximation, the vibrational transition energy is calculated to be 4318 cm-1 in the gas phase and 4242 cm-1 on the surface (without the Ag tip). The experimental value in the gas phase is 4161 cm-1 [48], and the harmonic equivalent, determined by fitting a harmonic potential to the bottom of the Morse-like potential, would correspond to 4401 cm-1 [49]. The slight discrepancy between the experimental and calculated frequencies are in line with the usual spurious softening of the PBE functional. Well-known scaling factors can be applied to bring the harmonic numbers closer to the anharmonic ones for this simple molecule. The calculations indicate a red-shift of 75 cm-1 upon the adsorption, which is larger but in reasonable agreement with the experimental TERS vibrational red-shift of H2/Ag(111). A simulated LT-TERS spectrum [46] is shown in Fig. S9 [39], which also agrees well with the experiment.
The TERS intensity is strongly correlated with the localized surface plasmonic resonance (LSPR) of the tip–sample junction. In general, there are two types of enhancement mechanisms effects in TERS [30, 38]: electromagnetic enhancement (originating from the LSPR of the junction with a plasmonic tip acting as an antenna for both incident light and Raman scattering) and chemical enhancement [originating from the interactions between the target molecule and the tip/surface]. In physisorption systems, the electromagnetic enhancement is expected to be dominant. We examined the excitation wavelength dependence (Fig. S10), verifying the contribution of LSPR to the TERS intensity; when the scattering light wavelength matches with the resonance range of the LSPR, strong TERS peak was observed.
The electromagnetic enhancement strongly depends on the cavity size. Figure 2(a) shows the gap-distance dependent TERS spectra recorded for D2/Ag(111). When the tip approaches the surface, the TERS intensities of the rotational and vibrational modes increase exponentially [Fig. 2(b)]. A broad background with a maximum at 590 nm is attributed to electronic Raman scattering which represents the spectral response of the LSPR in the junction [50]. The gap-distance dependent intensities of both rotational and vibrational modes are in good agreement with that of the plasmonic feature [Fig. 2(b)], indicating the dominant contribution of the gap-mode plasmon to the TERS enhancement.
The gap-distance dependence of the TERS intensities is well reproduced by simulating the electronic field distribution in the Ag-tip–Ag-surface junction using the finite element method (FEM) (see SM [39] for details). In the model, we assume a nanotip (0.5-nm radius of curvature) attached to a blunt tip body [51]. The black curve in Fig. 2(c) shows the field enhancement sampling at the position of a D2 molecule located under the Ag tip, indicating an exponential enhancement as the tip approaches. We confirmed the critical role of this sub-nanometric structure in TERS; as shown by the orange curve in Fig. 2(c), the tip without the nanotip exhibits a more moderate increase of the field enhancement as the tip approaches, which does not fit the experimental data [Fig. 2(b)]. The red curve in Fig. 2(c) shows the field enhancement at a possible position of the nearest neighboring molecule [Fig. 2(c), inset], indicating much lower enhancement. Therefore, TERS will have “nearly” single-molecule sensitivity, even for the physisorption system.
The rotational/vibrational spectroscopy is sensitive to the interaction between a target molecule and a local environment. We examine how the picocavity influences the rotational/vibrational state of H2 and D2. Figures 3(a)–3(d) show the gap-distance dependence of high-resolution TERS spectra focusing on the rotational/vibrational modes of H2/D2; notably, only the vibrational frequency of H2 exhibits a significant change [Fig. 3(c)]. Figure 3(e) shows the TERS spectra at three tip heights marked by the arrows in Fig. 3(c), indicating the red-shift and peak broadening as the gap distance decreases. Peak-fitting analysis for H2 [red dotted curve in Fig. 3(f)] indicates a corresponding peak shift of 34 cm-1 with a 200 pm displacement of the tip, while it is only 3 cm-1 for D2 (blue dotted curve). On the other hand, the rotational modes for both H2 and D2 [red and blue solid curves in Fig. 3(f), respectively] are almost unchanged, in good agreement with the calculated interatomic distance independent of the tip height as described above.
The significant (almost negligible) change of the vibrational mode of H2 (D2) can be explained by the anharmonicity of the potential energy surface (PES) of the molecule in the picocavity. The mode softening (red-shift and broadening) is associated with the PES of the molecule–surface distance, which is modified by the gap distance through the vdW interaction with the tip. Red-/Blue-shift of the intramolecular stretching mode through the vdW interaction has been reported by far-field Raman spectroscopy of hydrogen molecules physisorbed on carbon nanotubes [52]. The PES modification by the tip proximity was already reported for Au-tip–H2–Au-surface junction [53]; however, we clarified that the intermolecular interactions also play an important role in the PES shape (Figs. S4–S6 [39]). Since H2 has a higher zero-point energy (ZPE) than D2, their vibrational frequencies can be more affected by the PES modification. Anharmonicity may also modify the lifetime of molecular vibrations [54, 55], which could lead to the peak broadening. However, the influence of neighboring molecules may not be negligible [Fig. 2(c)], which would contribute to a shoulder at a lower frequency side of the vibrational peak [Fig. 3(e)].
To gain further insights into the observed red-shift, we calculated the PES experienced by the H2/D2 molecules directly under the tip at several gap distances . The PES is obtained by calculating the total energies by moving the molecule at several molecule–surface distances in the gap while the other atoms are fixed after the DFT optimization. The calculated PES exhibits a transition from a Morse-like potential [Fig. 3(h)] to an asymmetric double-well potential [Fig. 3(j)] as the gap distance decreases (see also Figs. S4 and S5). The quantum-corrected position for the H2/D2 molecule directly under the tip is obtained by solving the one-dimensional Schrödinger equation in the sinc-function discrete-variable representation (see detailed method, Fig. S6, and Table S1 in SM [39]). In parallel, we calculated the vibrational frequency of H2/D2 at each and with the i-PI code [56]. We found a hallmark of vibrational coupling, i.e., -dependent vibrational frequencies (Figs. S7 and S8). Figure 3(g) [3(i)] shows at = 0.94 [0.74] nm, where the stretching frequency is calculated to be 4247 [4230] cm-1. At the far [close] tip distance, is calculated to be 320 [341] pm, while = 316 [333] pm, as shown in Fig. 3(h) [3(i)]. This indicates that D2 is closer to the surface and farther from the tip than H2, due to the ZPE difference, and that both H2 and D2 molecules are slightly attracted toward the tip by the change from 0.94 to 0.74 nm. By the tip proximity, we predict 18 cm-1 red-shift for H2, but only 9 cm-1 red-shift for D2. This result is qualitatively consistent with the experiment, but under(over)estimates the red-shift for H2 (D2). The absolute gap distance may be one of the possible reasons for this discrepancy; the gap distance in the experiment will be slightly different between H2 and D2 even with the same STM setpoint because of different tip-apex geometries. Finally, because the calculations regard the quantum-mechanical electronic ground-state only, we cannot rule out that non-adiabatic effects in the junction could further influence the observed isotopic differences in the TERS spectra. Such an analysis will be the subject of future work.
With LT-TERS, we observed the rotational/vibrational transitions of single hydrogen molecules through the interaction with the picocavity field. Our results suggest that LT-TERS is applicable to weakly adsorbed molecules, offering deeper insights into the local structures, reactions, and dynamics of adsorbates, such as chemical reactions at active sites [57] and surface diffusion (spillover) in catalytic processes [58], at the single-molecule level.
Acknowledgements.
We thank Adnan Hammud for the Ag-tip fabrication. We thank Jun Yoshinobu and Heiko Appel for helpful discussions. T.S. acknowledges the support of JST FOREST Grant No. JPMJFR221U. T.K. acknowledges the support of JST FOREST Grant No. JPMJFR201J and JSPS KAKENHI Grant No. JP19K24684.References
- Ockwig and Nenoff [2007] N. W. Ockwig and T. M. Nenoff, Membranes for hydrogen separation, Chem. Rev. 107, 4078 (2007).
- Zhang et al. [2019] L. Zhang, M. Zhou, A. Wang, and T. Zhang, Selective hydrogenation over supported metal catalysts: from nanoparticles to single atoms, Chem. Rev. 120, 683 (2019).
- Li et al. [2020] X. Li, X. Ma, J. Zhang, E. Akiyama, Y. Wang, and X. Song, Review of hydrogen embrittlement in metals: hydrogen diffusion, hydrogen characterization, hydrogen embrittlement mechanism and prevention, Acta Metallurgica Sinica 33, 759 (2020).
- Fukutani and Sugimoto [2013] K. Fukutani and T. Sugimoto, Physisorption and ortho–para conversion of molecular hydrogen on solid surfaces, Prog. Surf. Sci 88, 279 (2013).
- Christmann [1988] K. Christmann, Interaction of hydrogen with solid surfaces, Surf. Sci. Rep. 9, 1 (1988).
- Vidali et al. [1991] G. Vidali, G. Ihm, H.-Y. Kim, and M. W. Cole, Potentials of physical adsorption, Surf. Sci. Rep. 12, 135 (1991).
- Bruch et al. [2007] L. Bruch, R. Diehl, and J. Venables, Progress in the measurement and modeling of physisorbed layers, Rev. Mod. Phys. 79, 1381 (2007).
- Avouris et al. [1982] P. Avouris, D. Schmeisser, and J. Demuth, Observation of rotational excitations of H2 adsorbed on Ag surfaces, Phys. Rev. Lett. 48, 199 (1982).
- Andersson and Harris [1982] S. Andersson and J. Harris, Observation of rotational transitions for H2, D2, and HD adsorbed on Cu(100), Phys. Rev. Lett. 48, 545 (1982).
- Yu et al. [1983] C.-f. Yu, K. B. Whaley, C. S. Hogg, and S. J. Sibener, Selective adsorption resonances in the scattering of -H2, -H2, -D2 and -D2 from Ag(111), Phys. Rev. Lett. 51, 2210 (1983).
- Anger et al. [1989] G. Anger, A. Winkler, and K. Rendulic, Adsorption and desorption kinetics in the systems H2/Cu(111), H2/Cu(110) and H2/Cu(100), Surf. Sci. 220, 1 (1989).
- Fukutani et al. [2003] K. Fukutani, K. Yoshida, M. Wilde, W. Diño, M. Matsumoto, and T. Okano, Photostimulated desorption and ortho-para conversion of H2 on Ag surfaces, Phys. Rev. Lett. 90, 096103 (2003).
- Niki et al. [2008] K. Niki, T. Kawauchi, M. Matsumoto, K. Fukutani, and T. Okano, Mechanism of the ortho-para conversion of hydrogen on Ag surfaces, Phys. Rev. B 77, 201404(R) (2008).
- Gupta et al. [2005] J. Gupta, C. Lutz, A. Heinrich, and D. Eigler, Strongly coverage-dependent excitations of adsorbed molecular hydrogen, Phys. Rev. B 71, 115416 (2005).
- Temirov et al. [2008] R. Temirov, S. Soubatch, O. Neucheva, A. C. Lassise, and F. S. Tautz, A novel method achieving ultra-high geometrical resolution in scanning tunnelling microscopy, New J. Phys. 10, 053012 (2008).
- Lotze et al. [2012] C. Lotze, M. Corso, K. J. Franke, F. von Oppen, and J. I. Pascual, Driving a macroscopic oscillator with the stochastic motion of a hydrogen molecule, Science 338, 779 (2012).
- Li et al. [2013] S. Li, A. Yu, F. Toledo, Z. Han, H. Wang, H. He, R. Wu, and W. Ho, Rotational and vibrational excitations of a hydrogen molecule trapped within a nanocavity of tunable dimension, Phys. Rev. Lett. 111, 146102 (2013).
- Natterer et al. [2013] F. D. Natterer, F. Patthey, and H. Brune, Distinction of nuclear spin states with the scanning tunneling microscope, Phys. Rev. Lett. 111, 175303 (2013).
- Natterer et al. [2014] F. D. Natterer, F. Patthey, and H. Brune, Resonant-enhanced spectroscopy of molecular rotations with a scanning tunneling microscope, ACS Nano 8, 7099 (2014).
- Therrien et al. [2015] A. J. Therrien, A. Pronschinske, C. Murphy, E. A. Lewis, M. L. Liriano, M. D. Marcinkowski, and E. C. H. Sykes, Collective effects in physisorbed molecular hydrogen on Ni/Au(111), Phys. Rev. B 92, 161407(R) (2015).
- Liu et al. [2018] S. Liu, A. Shiotari, D. Baugh, M. Wolf, and T. Kumagai, Enhanced resolution imaging of ultrathin ZnO layers on Ag(111) by multiple hydrogen molecules in a scanning tunneling microscope junction, Phys. Rev. B 97, 195417 (2018).
- Wang et al. [2022] L. Wang, Y. Xia, and W. Ho, Atomic-scale quantum sensing based on the ultrafast coherence of an H2 molecule in an STM cavity, Science 376, 401 (2022).
- Wang et al. [2023] L. Wang, D. Bai, Y. Xia, and W. Ho, Electrical manipulation of quantum coherence in a two-level molecular system, Phys. Rev. Lett. 130, 096201 (2023).
- Thijssen et al. [2006] W. Thijssen, D. Djukic, A. Otte, R. Bremmer, and J. van Ruitenbeek, Vibrationally induced two-level systems in single-molecule junctions, Phys. Rev. Lett. 97, 226806 (2006).
- Halbritter et al. [2008] A. Halbritter, P. Makk, S. Csonka, and G. Mihály, Huge negative differential conductance in Au–H2 molecular nanojunctions, Phys. Rev. B 77, 075402 (2008).
- Li et al. [2015] Y. Li, S. Kaneko, S. Fujii, and M. Kiguchi, Symmetry of single hydrogen molecular junction with Au, Ag, and Cu electrodes, J. Phys. Chem. C 119, 19143 (2015).
- Trouwborst et al. [2009] M. Trouwborst, E. Huisman, S. van der Molen, and B. van Wees, Bistable hysteresis and resistance switching in hydrogen-gold junctions, Phys. Rev. B 80, 081407(R) (2009).
- Ding et al. [2016] S.-Y. Ding, J. Yi, J.-F. Li, B. Ren, D.-Y. Wu, R. Panneerselvam, and Z.-Q. Tian, Nanostructure-based plasmon-enhanced Raman spectroscopy for surface analysis of materials, Nat. Rev. Mater. 1, 16021 (2016).
- Wang et al. [2020] X. Wang, S.-C. Huang, S. Hu, S. Yan, and B. Ren, Fundamental understanding and applications of plasmon-enhanced Raman spectroscopy, Nat. Rev. Phys. 2, 253 (2020).
- Itoh et al. [2023] T. I. Itoh, M. Procházka, Z.-C. Dong, W. Ji, Y. S. Yamamoto, Y. Zhang, and Y. Ozaki, Toward a new era of SERS and TERS at the nanometer scale: From fundamentals to innovative applications, Chem. Rev. 123, 1552 (2023).
- Höppener et al. [2024] C. Höppener, J. Aizpurua, H. Chen, S. Gräfe, A. Jorio, S. Kupfer, Z. Zhang, and V. Deckert, Tip-enhanced Raman scattering, Nat. Rev. Methods Primers 4, 47 (2024).
- Benz et al. [2016] F. Benz, M. K. Schmidt, A. Dreismann, R. Chikkaraddy, Y. Zhang, A. Demetriadou, C. Carnegie, H. Ohadi, B. De Nijs, R. Esteban, J. Aizpurua, and J. J. Baumberg, Single-molecule optomechanics in “picocavities”, Science 354, 726 (2016).
- Urbieta et al. [2018] M. Urbieta, M. Barbry, Y. Zhang, P. Koval, D. Sánchez-Portal, N. Zabala, and J. Aizpurua, Atomic-scale lightning rod effect in plasmonic picocavities: A classical view to a quantum effect, ACS Nano 12, 585 (2018).
- Zhang et al. [2013] R. Zhang, Y. Zhang, Z. Dong, S. Jiang, C. Zhang, L. Chen, L. Zhang, Y. Liao, J. Aizpurua, Y. Luo, J. L. Yang, and J. G. Hou, Chemical mapping of a single molecule by plasmon-enhanced Raman scattering, Nature 498, 82 (2013).
- Lee et al. [2019] J. Lee, K. T. Crampton, N. Tallarida, and V. A. Apkarian, Visualizing vibrational normal modes of a single molecule with atomically confined light, Nature 568, 78 (2019).
- Liu et al. [2023] S. Liu, F. P. Bonafe, H. Appel, A. Rubio, M. Wolf, and T. Kumagai, Inelastic light scattering in the vicinity of a single-atom quantum point contact in a plasmonic picocavity, ACS Nano 17, 10172 (2023).
- Xu et al. [2021] J. Xu, X. Zhu, S. Tan, Y. Zhang, B. Li, Y. Tian, H. Shan, X. Cui, A. Zhao, Z. Dong, J. Yang, Y. Luo, B. Wang, and J. G. Hou, Determining structural and chemical heterogeneities of surface species at the single-bond limit, Science 371, 818 (2021).
- Chen and Jensen [2023] R. Chen and L. Jensen, Interpreting chemical enhancements of surface-enhanced Raman scattering, Chem. Phys. Rev. 4 (2023).
- [39] See Supplemental Material for detailed experimental and theoretical methods, additional figures, and tables, which includes Refs. [59, 40, 16, 18, 19, 41, 42, 43, 60, 44, 56, 53, 61, 62, 46, 51, 63, 64, 65, 4, 48, 8, 66, 12, 13, 9, 67].
- Liu et al. [2020] S. Liu, B. Cirera, Y. Sun, I. Hamada, M. Müller, A. Hammud, M. Wolf, and T. Kumagai, Dramatic enhancement of tip-enhanced Raman scattering mediated by atomic point contact formation, Nano Lett. 20, 5879 (2020).
- Blum et al. [2009] V. Blum, R. Gehrke, F. Hanke, P. Havu, V. Havu, X. Ren, K. Reuter, and M. Scheffler, Ab initio molecular simulations with numeric atom-centered orbitals, Comput. Phys. Commun. 180, 2175 (2009).
- Gavini et al. [2023] V. Gavini, S. Baroni, V. Blum, D. R. Bowler, A. Buccheri, J. R. Chelikowsky, S. Das, W. Dawson, P. Delugas, M. Dogan, C. Draxl, G. Galli, L. Genovese, P. Giannozzi, M. Giantomassi, X. Gonze, M. Govoni, F. Gygi, A. Gulans, J. M. Herbert, S. Kokott, T. D. Kühne, K.-H. Liou, T. Miyazaki, P. Motamarri, A. Nakata, J. E. Pask, C. Plessl, L. E. Ratcliff, R. M. Richard, M. Rossi, R. Schade, M. Scheffler, O. Schütt, P. Suryanarayana, M. Torrent, L. Truflandier, T. L. Windus, Q. Xu, V. W.-Z. Yu, and D. Perez, Roadmap on electronic structure codes in the exascale era, Model. Simul. Mat. Sci. Eng. 31, 063301 (2023).
- Ruiz et al. [2012] V. G. Ruiz, W. Liu, E. Zojer, M. Scheffler, and A. Tkatchenko, Density-functional theory with screened van der Waals interactions for the modeling of hybrid inorganic-organic systems, Phys. Rev. Lett. 108, 146103 (2012).
- Kunisada and Kasai [2015] Y. Kunisada and H. Kasai, Hindered rotational physisorption states of H2 on Ag(111) surfaces, Phys. Chem. Chem. Phys. 17, 19625 (2015).
- Smeets and Kroes [2021] E. W. Smeets and G.-J. Kroes, Designing new SRP density functionals including non-local vdW-DF2 correlation for H2 + Cu(111) and their transferability to H2 + Ag(111), Au(111) and Pt(111), Phys. Chem. Chem. Phys. 23, 7875 (2021).
- Litman et al. [2023] Y. Litman, F. P. Bonafé, A. Akkoush, H. Appel, and M. Rossi, First-principles simulations of tip enhanced Raman scattering reveal active role of substrate on high-resolution images, J. Phys. Chem. Lette. 14, 6850 (2023).
- Pockrand and Otto [1981] I. Pockrand and A. Otto, Surface enhanced raman scattering (SERS): Annealing the silver substrate, Solid State Commun. 38, 1159 (1981).
- Veirs and Rosenblatt [1987] D. K. Veirs and G. M. Rosenblatt, Raman line positions in molecular hydrogen: H2, HD, HT, D2, DT, and T2, J. Mol. Spectrosc. 121, 401 (1987).
- Huber and Herzberg [1979] K. P. Huber and G. Herzberg, Molecular spectra and molecular structure: Constants of diatomic molecules (Van Nostrand, New York, 1979) pp. 716.
- Kamimura et al. [2022] R. Kamimura, T. Kondo, K. Motobayashi, and K. Ikeda, Surface-enhanced electronic Raman scattering at various metal surfaces, Phys. Status Solidi B 259, 2100589 (2022).
- Liu et al. [2019] S. Liu, M. Müller, Y. Sun, I. Hamada, A. Hammud, M. Wolf, and T. Kumagai, Resolving the correlation between tip-enhanced resonance raman scattering and local electronic states with 1 nm resolution, Nano Lett. 19, 5725 (2019).
- Williams et al. [2002] K. A. Williams, B. K. Pradhan, P. C. Eklund, M. K. Kostov, and M. W. Cole, Raman spectroscopic investigation of H2, HD, and D2 physisorption on ropes of single-walled, carbon nanotubes, Phys. Rev. Lett. 88, 165502 (2002).
- Sugimoto et al. [2017] T. Sugimoto, Y. Kunisada, and K. Fukutani, Inelastic electron tunneling mediated by a molecular quantum rotator, Phys. Rev. B 96, 241409(R) (2017).
- Persson and Ryberg [1989] B. Persson and R. Ryberg, Vibrational line shapes of low-frequency adsorbate modes: CO on Pt(111), Phys. Rev. B 40, 10273 (1989).
- Ueba [1986] H. Ueba, Vibrational lineshapes of adsorbates on solid surfaces, Prog. Surf. Sci. 22, 181 (1986).
- Kapil et al. [2019] V. Kapil, M. Rossi, O. Marsalek, R. Petraglia, Y. Litman, T. Spura, B. Cheng, A. Cuzzocrea, R. H. Meiß ner, D. M. Wilkins, B. A. Helfrecht, P. Juda, S. P. Bienvenue, W. Fang, J. Kessler, I. Poltavsky, S. Vandenbrande, J. Wieme, C. Corminboeuf, T. D. Kühne, D. E. Manolopoulos, T. E. Markland, J. O. Richardson, A. Tkatchenko, G. A. Tribello, V. V. Speybroeck, and M. Ceriotti, i-PI 2.0: A universal force engine for advanced molecular simulations, Comput. Phys. Commun. 236, 214 (2019).
- Cai et al. [2023] Z.-F. Cai, N. Kumar, and R. Zenobi, Probing on-surface chemistry at the nanoscale using tip-enhanced Raman spectroscopy, CCS Chem. 5, 55 (2023).
- Yang et al. [2024] J.-L. Yang, Y.-W. Zhou, M.-Y. Jia, Q. Liu, and Z.-F. Cai, Insights into heterogeneous catalysis by tip-enhanced Raman spectroscopy, ChemCatChem 16, e202400370 (2024).
- Summers [1969] R. L. Summers, Empirical observations on the sensitivity of hot cathode ionization type vacuum gages, NASA Technical Report NASA-TN-D-5285 (1969).
- Hermann and Tkatchenko [2020] J. Hermann and A. Tkatchenko, Density functional model for van der Waals interactions: Unifying many-body atomic approaches with nonlocal functionals, Phys. Rev. Lett. 124, 146401 (2020).
- Schwartz [1985] C. Schwartz, High-accuracy approximation techniques for analytic functions, J. Math. Phys. 26, 411 (1985).
- Colbert and Miller [1992] D. T. Colbert and W. H. Miller, A novel discrete variable representation for quantum mechanical reactive scattering via the -matrix Kohn method, J. Chem. Phys. 96, 1982 (1992).
- Jiang et al. [2016] Y. Jiang, S. Pillai, and M. A. Green, Realistic silver optical constants for plasmonics, Sci. Rep. 6, 30605 (2016).
- Pettinger [2010] B. Pettinger, Single-molecule surface-and tip-enhanced Raman spectroscopy, Mol. Phys. 108, 2039 (2010).
- Wolniewicz [1983] L. Wolniewicz, The state vibration-rotational energies of the H2, HD, and D2 molecules, J. Chem. Phys. 78, 6173 (1983).
- Sakurai et al. [1988] M. Sakurai, T. Okano, and Y. Tuzi, Ortho-para conversion of -H2 physisorbed on Ag(111) near two-dimensional condensation conditions, Appl. Surf. Sci. 33, 245 (1988).
- Teal and MacWood [1935] G. K. Teal and G. E. MacWood, The Raman spectra of the isotopic molecules H2, HD, and D2, J. Chem. Phys. 3, 760 (1935).
Supplemental Material: Picocavity-enhanced Raman Spectroscopy of Physisorbed H2 and D2 Molecules
I I. Methods
I.1 A. STM-TERS Experiments
The experiments were performed in an ultrahigh vacuum (UHV) chamber equipped with an STM-based TERS machine (modified UNISOKU USM-1400) at a sample temperature of 10 K. A single-crystalline Ag(111) surface (MaTeck) was cleaned by Ar-sputtering-and-annealing cleaning cycles in the chamber. H2 (99.999%, Westfalen AG) or D2 (99.7%, Air Liquide) gas was dosed into the STM unit cooled at 10 K via a variable leak valve. The gas exposure (Langmuir; L) is estimated by , where [ Torr], [s], and denote the displayed pressure value of the gauge (corrected by N2), the exposure duration, and the correction factor for the gauge. A Bayard-Alpert ionization gauge is located outside the cryogen-cooled STM unit in the chamber and (0.388) is used for H2 (D2) [59]. We used a chemically-etched Ag tip that was sharpened by focused-ion-beam milling as described elsewhere [40]. The bias voltage was applied to the sample while the tip was grounded. The STM images were obtained in the constant current mode at = 50 pA. The cleanness of the cleaned surface and tip was confirmed by STM images [Fig. S1(a)]. We also confirmed that a TERS spectrum of the Ag–Ag junction before hydrogen dosing had no molecule-derived peaks (Fig. S2, at 0 L)
In the STM junction, -polarised visible laser (532-nm solid-state or 633-nm HeNe) was incident through a fused silica window and a Ag-coated parabolic mirror (numerical aperture of 0.6). The incident laser power is 5.9 (9.7) mW at = 532 (633) nm. The parabolic mirror is mounted on a five-axis piezoelectric-motor stage in the LT-STM unit and the laser is focused at the tip apex by adjusting the position and angle of the parabolic mirror. The scattered light corrected by the parabolic mirror was directed to a spectrometer (AndorShamrock 303i) outside the UHV chamber via a beam splitter and a longpass filter specific for each excitation wavelength. For the TERS and STM-induced luminescence (STML) spectra, we used the gratings 150 (for Fig. S3), 600 (Figs. 1, 2, and S2), and 1200 line/mm (Fig. 3). A sparser grating allows the acquisition of wider-wavelength range spectra, as shown in Fig. S3 (see Sec. II-B), but Raman peaks become broad and asymmetric. Thus, for the high-frequency-resolution measurements shown in Fig. 3, the densest grating was used.
I.2 B. DFT Calculations and Molecular Dynamics
DFT calculations were performed with the FHI-aims program package [41, 42]. We used intermediate defaults for basis sets, radial grids and cutoff potential parameters and light defaults for the Hartree multipole expansion and angular grids. This combination proved efficient and reliable against benchmarks to tight settings. We simulated an orthogonal Ag(111) surface cell oriented perpendicular to the axis, containing four layers. The cell parameters used were Å, Å and Å. We employed a -point grid and a dipole correction was employed to maximally decouple periodic images along the direction. We used the PBE exchange-correlation functional and included vdW interactions through the pairwise screened TS-vdWsurf method of Ref. [43] in all simulations reported in the main text. Figure S4 shows a comparison of binding energies with the many-body NL-MBD dispersion correction of Ref. [60], indicating a weaker binding energy by only 5 meV and a binding distance 0.1 Å closer to the surface. Structures were optimized by allowing only hydrogen molecules and the top layer of Ag atoms of the surface to relax. All the other atoms were kept fixed at their bulk (or tip) positions. When we performed simulations including a model tip structure, this was modeled by a Ag19 cluster at varying distances from the surface. Although this tip size would not be sufficient to capture a proper plasmonic response of the tip, the tip structure can capture the local interactions close to the tip apex [46, 36].
We inspected binding sites and binding energies of a single adsorbed H2 molecule, finding all values in good agreement with previous studies [44]. The binding energy was defined as the difference between the total energy when the molecule was 10 Å away from the surface and the minimum total energy of the adsorbed molecule. The blue curve in Fig. S4 shows the binding curve of a single H2 molecule adsorbed on the fcc-hollow site of the Ag(111) surface, indicating an equilibrium H2–Ag(111) bonding length of 3.09 Å and a binding energy of 32 meV. The model we used in most simulations contained 11 molecules in the unit cell (see rationale below). For this molecular concentration, we evaluated the intermolecular interaction by comparing the binding curves (orange and green curves in Fig. S4) with that of the single-molecule case (blue); the molecule–molecule interaction in our models amounts to 55 meV/molecule.
Performing vibrational analysis on the optimized single molecule adsorbed on the surface showed that negative frequencies related to rotational and translational modes of the molecule were always present. This led us to consider higher coverages to model local stable clustering of the molecules on the surface. We considered a full monolayer (1 ML), where all fcc-hollow sites of the unit cell were occupied by H2. With this setup we performed Born-Oppenheimer molecular dynamics (BOMD) and path-integral molecular dynamics (PIMD) at LT ( = 20 K) with the i-PI code [56] connected to FHI-aims. PIMD simulations used 12 beads in conjunction with a normal-mode colored-noise thermostat with 8 auxiliary variables and . This number of beads is not fully converged at this temperature even with the colored-noise quantum thermostats, but the purpose of these simulations was exploratory.
While the full monolayer was stable during more than 10 ps on the classical trajectories, in the quantum trajectory 5 molecules readily desorbed (within the first picosecond) from the substrate. For the quantum case we validated this observation by simulating trajectories with a parameterized Lennard-Jones potential representing the Ag(111) surface. We could calculate the radius of gyration of H atoms from the PIMD simulations, belonging to the 3 molecules closest to the tip at any point in the simulation. For the closest tip distance considered ( = 7.4 Å), these were 0.33 Å along and and 0.30 Å along ( = 20 K), indicating a small anisotropy. Converting these numbers to an effective thermal de Broglie wavelength and applying a inverse-square-root scaling for the temperature and mass, we would obtain Å and Å at 10 K. We note that, depending on the density of H2 molecules on the surface, is about the length of the inter-particle spacing at 10 K and signatures of bosonic exchange could begin to be observed. These features are disregarded in the current simulations.
Figure S5 shows the energy profiles of a H2 molecule with varying distance to the Ag(111) surface at different tip-surface distances. We only varied the position of the molecule closest to the tip apex, while the others were kept at equilibrium [see also Figs. 3(g) and 3(i) in the main text]. The profile shapes are similar to an asymmetric double-well potential at closer tip–surface distances, while they assume a shape closer to a Morse potential at larger distances, which asymptotically approaches the profile of a H2 molecule on Ag(111) without the Ag tip (yellow curve in Fig. S5). The asymmetry of the double-well is strongly impacted by the presence of other molecules on the surface (molecule–molecule interactions). Figure S6 shows a comparison with the limiting case of a single molecule in the unit cell at a tip–surface distance of 9.4 Å. In that case (orange curve), if the molecule can overcome the barrier, it will have a higher probability of being found closer to the tip, which was predicted in a Au-tip–single-H2–Au-surface junction [53].
I.3 C. Exact DVR Models and Vibrational Analysis
For the potentials shown in Fig. S5, we solved the one-dimensional Schrödinger equation exactly in the sinc-function discrete variable representation (DVR) [61, 62]. We set the mass of the hydrogen atom to and that of deuterium to . In all cases, calculations were performed on a regularly spaced grid of 100 points, spanning values of from to with a potential-energy cut-off . The potential energies used in the DVR were obtained from a cubic spline interpolation of the DFT-calculated values in Sec. I-B. The resulting ZPEs and average ground-state displacements of the molecular center of mass from equilibrium,
are given in Table S1, as well as those for D2. Here, is the molecular height considering the quantum mechanical ZPE contributions at the PES (i.e., the “quantum-corrected” molecular heights), and are the underlying ground-state wavefunctions, which are plotted in Fig. S7 to further illustrate the influence of tip–surface separation on the ZPEs and the equilibrium displacements.
no tip | ||||||
In parallel we conducted harmonic vibrational analysis with the i-PI code, varying the height of the H2/D2 molecule directly under the tip apex and varying the tip–sample distance [see also Figs. 3(g) and 3(i)]. Figures S9 and S9 show the dependence of the H–H and D–D stretching vibration on the mode corresponding to the vertical hindered translation of the adsorbed molecule, i.e., H2–Ag(111) and D2–Ag(111), respectively. At each , there is a red-shift of the interatomic stretching frequency with approaching tip, which follows a mass scaling between H2 and D2. As described above, with the DVR solutions, we obtain the the quantum-corrected molecule–surface distance . Based on the -dependent vibrational frequency curves, we obtain the vibrational frequency at , as shown by the blue (green) squares for H2 (D2) in Fig. S9 (S9). The plot indicates an anomalous isotopic red-shift for H2 (factor 2, instead of ), stemming from the two following factors: (1) strong anharmonic character of the hindered translation of the molecule, which is dominated by the vdW interactions with the tip and the surface (see Fig. S5), and (2) the coupling of this mode with the vibrational stretching mode.
We stress that different heights of the tip and different concentrations of the molecules can exacerbate this effect. The former because the red-shift is more pronounced at closer tip distances than at larger tip distances. The latter because weaker binding energies to the surface at lower concentrations result in softer potentials that assume a double-well character at larger tip distances (see the orange curve in Fig. S6). This effect can result in a more pronounced negative slope of the dashed curves in Figs. S9 and S9. It is challenging to determine the local molecular concentration closer to the tip apex at any measurement and a deeper study of of the vibrational frequency variation could provide indirect hints to the observed concentration. Although we neglected the anharmonic character of the H–H stretching coordinate in this study, we expect this effect not to change the qualitative discussion and introduce only rigid red-shifts on the values reported here.
I.4 D. TERS Spectrum Simulation
We conducted TERS simulations according to the methodology proposed in Ref. [46]. We used Tip A discussed in the reference and calculated the anharmonic Raman spectrum based on a classical ab initio molecular dynamics (AIMD) trajectory (5 ps) of a molecular layer adsorbed on the surface and a tip apex located 4 Å above the molecular layer at = 20 K. We calculated the TERS intensity as , where was obtained by a local polarizability calculation at each geometry in the time-series given by the molecular dynamics. For this part of the calculation we only considered the free-standing molecular layer (no surface) at the positions adopted in the AIMD trajectory including the surface. We note that the method of Ref. [46] is unable to capture the electromagnetic plasmonic enhancement in its totality. We do not observe any noteworthy features in the spectrum and it overall agrees with the experimental spectrum. The red-shift of the rotational band in the simulation is explained by the lack of consideration of the nuclear spin, making calculations reproduce only classical rotations that are red-shifted if compared to the quantized transition of para-H2. This simulation had only an exploratory character and achieving statistical convergence through the use of parametrized machine-learned potentials, plus the inclusion of accurate plasmonic enhancements is beyond the scope of this paper.
I.5 E. FEM Simulations
To evaluate the electric field in the STM junction, we performed finite element method (FEM) simulations using COMSOL Multiphysics (version 6.2 with the Wave Optics Module). We used the three-dimentional model with a Ag tip and Ag sample plate described in Supporting Information of Ref. [51]. As needed, at the apex of the Ag tip (a length of 300 nm, an effective radius of 30 nm), a half-sphere protrusion with a radius of 0.5 nm is attached as a “nanotip.” A -polarized plane wave with of 532 nm and an electric field of 1 V/m is incident with an angle of 35∘ relative to the sample surface. The perfectly matched layers surrounding the volume are set to absorb all outgoing waves. The dielectric constant of the Ag objects is referenced from the literature [63]. The surface normal component of the electric field is sampled.
A Ag flat plate is placed beneath the tip apex with a gap distance . Here we define the origin of the gap distance ( = 0) as the tip height when the Ag-tip apex forms point contact with the Ag surface. As described above, the quantum-corrected molecular height is calculated to be approximately 0.32–0.33 nm for D2/Ag(111). In the definition, taking into account the atomic radius of Ag, the position of the molecule corresponds to 0.2 nm above the Ag surface plane in the FEM model. Although the calculated depends on the tip height [Figs. 3(h) and 3(j) in the main text], in the FEM model, we used the constant height for the sampling point at any gap distances. We confirmed that differences in the definition or the molecular position on the order of 0.1 nm do not result in qualitative differences in the field enhancement factor curves. To directly compare with the experimental TERS intensities, the electric field enhancement factor (Ref. [64]) is displayed in the plots of Fig. 2(c) in the main text. To the field enhancement of the nearest neighboring molecule, assuming the molecule is placed in the next adsorption site of Ag(111), = 0.29 nm is used [red triangle in the inset of Fig. 2(c) in the main text]. Thus, when the point at the surface just under the tip is defined as (, , ) = (0, 0, 0), the black and red curves shown in Fig. 2c of the main text represent the plots sampling at (, , ) = (0, 0, 0.20) nm and (0.29, 0, 0.20) nm, respectively.
The tip heights in the DFT calculations are comparable with the gap distances with the FEM simulations that reproduced the Raman enhancement [Fig. 2(c) in the main text; the gap distances of 550 to 850 pm in the simulation correspond to the tip height of 0.8 to 1.1 nm, considering the Ag-atom radius].
II II. Supplemental Results and Discussion
II.1 A. Nuclear Spin Isomers
H2 and D2 have two nuclear spin isomers, ortho and para. para-H2 (nuclear spin quantum number = 0) is more stable than ortho-H2 ( = 1) by 14 meV [65]. para- (ortho-) H2 has only even (odd) rotational quantum numbers . Although ortho-H2 is the predominant species in the gas phase at room temperature (a temperature equilibrium composition of 75%), the molecular adsorption on metals enhances the ortho-to-para conversion, leading to the dominance of the para species on the surfaces [4]. In the TERS spectra, we can detect only -H2 (Table S2); as a reference, in the gas phase, ortho-H2 have a Raman-active rotational mode [] at 587 cm-1 [65, 48]. This implies that we detected only the physisorbed species after ortho-para conversion on the surface. In the case of D2, ortho ( = 0, 2) is more stable than para ( = 1) by 7.4 meV and more dominant (67%) in the gas phase at room temperature. Only -D2 was detected in our TERS spectra, indicating again that para-D2 was converted on the surface.
Although the dominance of para-H2 in the TERS spectra is consistent with a previous HREELS study of H2 on a Ag film (0.5–10 L) [8], some previous reports detect ortho-H2 on Ag(111). A HREELS study on H2/Ag(111) [66] shows that rotational mode peaks of both para-H2 and ortho-H2 were detected under a higher H2 pressure ( Torr) while para-H2 was dominantly detected under a lower pressure ( Torr). Photostimulated desorption studies (30 L) [12, 13] shows that ortho-H2 and para-D2 are alive on Ag(111) on the order of hundred seconds and that the conversion of D2 is slower than H2.
To compare with the above-mentioned studies, we performed TERS measurements during exposing D2 to a clean Ag(111) surfaces, as shown in Fig. S2. The gas pressure was set to Torr, comparable with the EELS study [66]. The ortho-D2-derived peaks appeared at 16 L and then saturated at 40 L, while para-D2 was not detected. The weak intensity at low exposures may be due to molecular diffusion along the surface at low coverages and the weak contribution of neighboring molecules to the TERS signal [see Fig. 2(c) in the main text]. It should be noted that even if the gas pressure and exposure values are comparable, the molecular coverage varies depending on the chamber configuration. Nevertheless, we confirmed that ortho-H2 and para-D2 were not detected at any pressures and exposures we used (0.8–2 Torr; 102–103 Langmuir). The HREELS observation of the metastable isomer only at high exposures [66] may be attributed to be hydrogen multilayers, where the second layer does not interact with the Ag surface. In contrast, the narrow TERS gap for the LT-TERS measurement would suppress the formation of the multiple layers. Another possibility that the metastable isomer were not detected in TERS is that the LSPR in the gap is accelerating the para-to-ortho conversion of the molecule under the tip. Although far-field 532-nm irradiation did not affect the conversion rate [13], the strong electric field localized in the STM junction could interact with the nuclear spin of the adsorbate. This effect would be verified through theoretical calculations.
System | On Ag(111) | On Ag film | On Cu(001) | Gas phase | ||
Method | TERS (low res.) | TERS (high res.) | HREELSa | HREELSb | Ramanc | Ramand |
H2 | ||||||
rot | 351 | 348 | 395 | 363 | 354.5 | 354.365 |
vib | 4121 | 4134 | 4178 | 4178 | - | 4161.200 |
rot+vib | 4443 | - | 4533 | 4517 | - | 4497.848 |
D2 | ||||||
rot | 169 | 172 | - | 177 | 178.8 | 179.108 |
vib | 2967 | 2968 | - | 3000 | 2996.1 | 2993.6 |
rot+vib | 3133 | - | - | 3170 | 3166.3 | - |
II.2 B. Excitation Wavelength Dependence
To verify the contribution of LSPR to the TERS intensity, we examined the excitation wavelength () dependence. Figure S3(a) shows a STML spectrum recorded over a clean Ag(111) surface with a Ag tip. The LSPR results in a broad spectral feature in the visible range with a maximum at 560 nm. After recording the STML, we exposed the surface to the D2 gas, conducting TERS measurements with = 532 and 633 nm [Figs. S3(b) and S3(c), respectively]. The peak intensity of each mode significantly changes with , whereas the Raman shift value is unchanged. The vibrational intensity is weakened with = 633 nm, indicating the corresponding scattering light is not efficiently enhanced by the LSPR because the peak position ( = 779 nm) is located in the weak resonance range of the LSPR [Fig. S3(a)]. In contrast, both the rotational and vibrational peaks become intense with = 532 nm because they are more resonant with the LSPR.

### v2

@@
 Abstract
-We report on tip-enhanced Raman scattering (TERS) of H2 and D2 molecules physisorbed within a plasmonic picocavity at a cryogenic temperature (10 K).
-The intense Raman peaks resulting from the rotational and vibrational transitions are observed at sub-nanometer gap distances of the junction formed by a Ag tip and Ag(111) surface.
-We clarify that the predominant contribution of the electromagnetic field enhancement of the picocavity to the detection of a single hydrogen molecule.
-The gap-distance dependent TERS reveals not only the evolution of the picocavity field, but also the interaction between the molecule and tip/surface, which exhibit nontrivial isotope effects.
-A significant red-shift and peak broadening of the H–H stretching as the gap distance decreases, while the D–D stretching mode is unaffected.
-A combination of density functional theory and reduced-dimension models reveals that a distinct anharmonicity in the mode potential of H2 is one cause of the anomalous red-shift, whereas D2 has less anharmonicity due to the geometric isotope effect.
+We report on tip-enhanced Raman scattering (TERS) of H2 and D2 molecules physisorbed within a plasmonic picocavity at 10 K.
+The intense Raman peaks resulting from the rotational and vibrational transitions are observed at sub-nanometer gap distances of the junction formed by an Ag tip and Ag(111) surface, where a picocavity-enhanced field plays a crucial role.
+A significant red-shift of the H–H stretch frequency is found as the gap distance decreases, while the D–D stretch frequency is unaffected.
+Density functional theory, path-integral molecular dynamics, and quantum anharmonic vibrational energy calculations suggest that this unexpected isotope effect is explained by a different molecular density between H2 and D2 on the surface.
 The adsorption of hydrogen molecules on solid surfaces is the first step in permeation for fuel storage [1], catalytic hydrogenation [2], hydrogen embrittlement [3], and nuclear-spin isomer conversion [4].
 To understand and control the elementary processes in such physical and chemical phenomena, hydrogen adsorption has been intensively studied in surface science [5, 6, 7].
-However, only a few methods can be used to characterize physisorption systems, such as hydrogen on coinage-metal surfaces [8, 9, 10, 11, 12, 13], because cryogenic-temperature measurements are required due to the faint adsorption energy ( meV).
-Low-temperature scanning tunneling microscopy (LT-STM) has been employed to investigate the dynamics of hydrogen molecules weakly adsorbed on surfaces [14, 15, 16, 17, 18, 19, 20, 21, 22, 23].
-Several studies have assigned characteristic features in conductance spectra to the rotational mode of hydrogen in the tip–sample junction [17, 18, 19, 20].
-However, the interpretation of the conductance spectra remains controversial; the features within the corresponding energy range [30–60 meV (240–480 cm-1) for H2] could be attributed to configuration switches [14, 24, 25], to a molecule?substrate stretching mode [26], or to multiple phonon excitation [27].
-Additionally, LT-STM studies have not provided information on the interatomic stretching mode of the molecule.
-Therefore, an alternative approach is required to identify the rotational/vibrational state.
-Raman spectroscopy is a promising technique for characterizing hydrogen molecules on surfaces since the rotational and vibrational transitions of a homonuclear diatomic molecules are observable.
-However, the detection demands exceptional sensitivity due to the intrinsically small cross-section.
-Recent advancements in tip-enhanced and surface-enhanced Raman spectroscopy (TERS/SERS) have enabled the detection of single molecules [28, 29, 30, 31].
-Picocavities, atomic-scale protrusions present in plasmonic nanojunctions, play a crucial role in achieving extreme confinement of the electromagnetic field to Ångström scales [32, 33, 31], resulting in a significant enhancement of Raman scattering and ultrahigh spatial resolution [34, 35, 36, 37].
-So far, the exceptional sensitivity in TERS/SERS has been demonstrated primarily for organic molecules chemisorbed onto the substrate, where the enhancement benefits from chemical effects arising from the molecule–surface (and molecule–tip) interactions [38].
-Therefore, applying TERS/SERS to physisorbed systems is not straightforward, as it depends on the fundamental enhancement mechanisms of Raman scattering in plasmonic picocavities.
-Here we demonstrate that LT-TERS can observe single hydrogen molecules physisorbed onto an Ag(111) surface with an Ag tip [Fig.
-1(a), inset], identifying their rotational and vibrational states.
-Furthermore, precise gap-distance control enables us to investigate not only how the picocavity field contributes to the Raman scattering process but also how the subtle van der Waals (vdW) interactions between the molecule and the tip/surface affect the anharmonicity of the potential energy surfaces and the accompanying nuclear quantum effects.
-The LT-TERS experiments were performed in an ultra-high vacuum chamber, as described in the Supplemental Material (SM) [39].
-The 532-nm incident laser beam was linearly polarized along the tip axis and focused on the apex of a focused-ion-beam-sharpened Ag tip [40] using an in-situ parabolic mirror, and the scattering light was detected by an ex-situ spectrometer.
-We performed density functional theory (DFT) calculations [41, 42] on a model geometry setup.
-We considered H2/D2 on a Ag(111) (periodic) surface, employing the Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional and screened pairwise vdW interactions [43] (see SM [39] for details).
+Specifically, hydrogen, the lightest molecule, exhibits unique quantum effects that differ from those of other molecules when it is physisorbed on a surface through van der Waals (vdW) interactions [8, 9, 10, 11, 12, 13, 14, 15, 16, 17].
+However, only a few methods are available for characterizing such physisorption systems as the weak adsorption energy ( meV) necessitates measurements at cryogenic temperatures.
+Low-temperature scanning tunneling microscopy (LT-STM) has been employed as a local characterization technique to investigate the dynamics of hydrogen molecules weakly adsorbed on surfaces [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28].
+Several studies have observed characteristic features in conductance spectra, attributed to the rotational transition of molecular hydrogen in the STM junction [21, 22, 23, 24].
+However, the interpretation remains controversial; the features within the corresponding energy range [30–60 meV (240–480 cm-1) for H2] could be attributed to configuration switching [18, 29, 30], to a molecule–substrate stretching mode [31], or to multiple phonon excitation [32].
+Additionally, the stretching mode has never been observed in these studies.
+Raman spectroscopy is a promising technique for characterizing hydrogen molecules on surfaces, as it allows the observation of both rotational and vibrational transitions of homonuclear diatomic molecules.
+Although the detection demands exceptional sensitivity due to the intrinsically small cross-section, single-molecule detection can be achieved by tip-enhanced and surface-enhanced Raman spectroscopy (TERS/SERS) [33, 34, 35, 36, 37].
+Recent studies have established that picocavities, atomic-scale protrusions present in plasmonic nanojunctions, play a crucial role in achieving extreme confinement of the electromagnetic field to Ångström scales [38, 39, 37], resulting in a significant enhancement of Raman scattering [40, 41, 42] and ultrahigh spatial resolution [43, 44, 45, 46].
+Yet, the exceptional sensitivity in TERS/SERS has been demonstrated primarily for organic molecules chemisorbed onto the substrate, where the enhancement benefits from chemical effects arising from the molecule–surface (and molecule–tip) interactions [47].
+Therefore, applying TERS/SERS to physisorbed systems has remained challenging.
+Here we demonstrate that LT-TERS can detect hydrogen molecules physisorbed onto an Ag(111) surface with an Ag tip [Fig.
+1(a), inset], allowing the identification of their rotational and vibrational states.
+Furthermore, precise gap-distance control enables us to investigate not only the contribution of the picocavity field to the Raman scattering process but also the influence of tip proximity on the rotational/vibrational energies of the molecules through non-trivial nuclear quantum effects.
+Density functional theory (DFT), path-integral molecular dynamics (PIMD), and quantum vibrational energy calculations suggest that the tip–molecule–surface potential energy profiles are critically affected by the intermolecular and tip–molecular interactions, leading to unusual isotope effects on vibrational frequencies.
+The LT-TERS experiments were performed in an ultra-high vacuum chamber [see Sect.
+I-A of the Supplemental Material (SM) [48]].
+The 532-nm incident laser beam was linearly polarized along the tip axis and focused on the apex of a focused-ion-beam-sharpened Ag tip [40, 42] using an in-situ parabolic mirror, and the scattering light was detected by an ex-situ spectrometer.
+We conducted DFT calculations [49, 50] of H2/D2 on a Ag(111) (periodic) surface, employing the Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional and screened pairwise vdW interactions [51] (see SM Sect.
+I-B [48]).
 The preferred adsorption sites (fcc/hcp hollow) and binding energies (30 meV) were in very good agreement with previous studies of isolated molecules on the surface [52, 53].
-A single H2/D2 molecule proved not to be stable in our simulations, due to the extremely low diffusion barrier.
-Therefore, we considered a coverage of 0.69 monolayer (ML), motivated by the results of path-integral molecular dynamics simulations at LT, where a 1 ML coverage was also unstable (see Section I-B in SM [39]).
-In 0.69 ML, the intermolecular attractive interaction increases the binding energy by 55 meV (Fig.
-S3 [39]).
-Over the sample, we put a Ag19 cluster in a cleaved pyramidal geometry as a Ag tip (see Section I-B in SM [39] and Refs.
-[46, 36] for the tip structure details).
-H2 or D2 gas was dosed into the chamber where a clean Ag(111) surface was cooled at 10 K.
-Regardless of laser illumination, hydrogen molecules on the surface are invisible with LT-STM (Fig.
-S1 [39]).
+Unless specified otherwise, we considered a coverage of 0.69 monolayer (ML), motivated by the results of PIMD simulations at LT (see SM Sect.
+I-B [48]).
+At 0.69 ML coverage, the intermolecular attractive interaction increases the binding energy by 55 meV compared to the lowest coverage of 0.06 ML considered here (Fig.
+S4).
+Over the sample, we placed the Ag tip model shown in SM Sect.
+I-B [48] (see also Fig.
+S5), which was also previously discussed in Refs.
+[46, 54, 41].
+H2 or D2 gas was introduced into the chamber where a clean Ag(111) surface was cooled at 10 K (see SM Sect.
+I-A [48]).
+Regardless of laser illumination, no static hydrogen molecule on the surface was imaged with LT-STM (Fig.
+S1).
 In contrast, clear peaks were observed in TERS spectra under these conditions.
-Figure 1(a) displays a typical TERS spectrum of H2 (D2) on Ag(111), showing two intense peaks at 351 (169) and 4121 (2967) cm-1 together with a weak peak at 4443 (3133) cm-1.
-The three peaks in the TERS spectra can be assigned to rotational/vibrotational modes of hydrogen.
-Note that broad or asymmetric peak shapes are artifacts in the wide wavenumber range spectra (see Section I-A in SM [39]); higher resolution spectra will be shown later.
-The structure at 100 cm-1 common to the two spectra originates from Ag phonon modes of the tip [47, 36].
-Figure 1(b) depicts the energy diagram of rotational/vibrational levels and possible transitions under the Raman selection rule [the change of the vibrational (rotational) quantum numbers () from the initial to final states should be 0 or (0 or )].
+Figure 1(a) displays typical TERS spectra of H2 and D2 on Ag(111).
+For H2, intense peaks were observed at 351 and 4121 cm-1, along with a weak peak at 4443 cm-1.
+Similarly, for D2, intense peaks appeared at 169 and 2967 cm-1, with a weak peak at 3133 cm-1.
+The three peaks can be assigned to rotational/vibrotational modes of hydrogen.
+Note that broad or asymmetric peak shapes are artifacts in the wide wavenumber range spectra (see SM Sect.
+I-A [48]).
+The feature around 100 cm-1 common to both spectra originates from phonon modes of the Ag tip [41, 55], which is present even in the absence of hydrogen adsorption (Fig.
+S2).
+Figure 1(b) depicts the energy diagram of rotational/vibrational levels and allowed transitions under the Raman selection rule.
 The observed peaks, in order from lowest to highest, correspond to the lowest-energy rotational transition [; “rot” labeled in Fig.
 1(a)], the lowest-energy vibrational transition [; “vib”], and their combination [; “rot+vib”], respectively.
-The three modes were also detected in high-resolution electron energy loss spectroscopy (HREELS) of hydrogen on a polycrystalline Ag film [8] and a Cu(001) surface [9] at approximately 10–12 K; however, better energy resolution (10 cm-1) in TERS than that of the HREELS (28 cm-1 [9]) makes the red-shifts from gas-phase hydrogen more discernible (Table S2).
-We should also note that only the stable nuclear spin isomer of each isotope, i.e., para-H2 and ortho-D2 [ = even; Fig.
+The three modes were also observed in high-resolution electron energy loss spectroscopy (HREELS) of hydrogen on a polycrystalline Ag film [8] and a Cu(001) surface [9] at approximately 10–12 K; however, better energy resolution (10 cm-1) in TERS than that of the HREELS (28 cm-1 [9]) makes the red-shifts from gas-phase hydrogen more discernible (Table S4).
+We should also note that only the stable nuclear spin isomer of each isotope, i.e., para-H2 and ortho-D2 [Fig.
 1(b)], is detected in TERS (Fig.
 S2).
-On the surface, the stable isomer is dominant due to the ortho–para conversion [8, 4, 44] (see SM [39] for details).
+On the surface, the stable isomer is dominant due to the ortho–para conversion [8, 4, 52] (see SM Sect.
+II-A [48]).
+The rotational transition at 351 cm-1 for H2/Ag(111) [Fig.
+1(a)] is slightly red-shifted from 354 cm-1 in the gas phase [56].
 We calculate the H–H bonding distance to be 75.4 pm on Ag(111) (cf.
 75.1 pm in the gas phase).
-Considering a rigid-rotor approximation, the rotational transition energy of para-H2 is calculated to be of 353 cm-1 on the surface, compared to 356 cm-1 in the gas phase.
-This is in good agreement with the experimental TERS peak at 351 cm-1 (Fig.
-1a) slightly red-shifted from 354 cm-1 in the gas phase [48].
-Notably, the H–H bonding distance of H2 on Ag(111) varied by only 0.06 pm at the closest tip position with respect to that without any tip included in our simulations.
+Considering a rigid-rotor approximation, the rotational transition energy of para-H2 is calculated to be 353 cm-1 on the surface, compared to 356 cm-1 in the gas phase.
+This is in good agreement with the experiment.
+Notably, the H–H bonding distance of H2 on Ag(111) varied by only 0.06 pm at the closest tip position with respect to simulations without any tip.
 This suggests that the hydrogen molecules on the surface and in the STM junction behave as nearly free rotors.
-In the harmonic approximation, the vibrational transition energy is calculated to be 4318 cm-1 in the gas phase and 4242 cm-1 on the surface (without the Ag tip).
-The experimental value in the gas phase is 4161 cm-1 [48], and the harmonic equivalent, determined by fitting a harmonic potential to the bottom of the Morse-like potential, would correspond to 4401 cm-1 [49].
-The slight discrepancy between the experimental and calculated frequencies are in line with the usual spurious softening of the PBE functional.
-Well-known scaling factors can be applied to bring the harmonic numbers closer to the anharmonic ones for this simple molecule.
-The calculations indicate a red-shift of 75 cm-1 upon the adsorption, which is larger but in reasonable agreement with the experimental TERS vibrational red-shift of H2/Ag(111).
-A simulated LT-TERS spectrum [46] is shown in Fig.
-S9 [39], which also agrees well with the experiment.
+In the TERS spectrum [Fig.
+1(a)], the vibrational frequency of 4121 cm-1 exhibits a red-shifted relative to the gas-phase value of 4161 cm-1 [56].
+The vibrational transition energy is calculated to be 4088 cm-1 on the surface (without the Ag tip), after scaling the calculations performed in the harmonic approximation (see SM Sect.
+I-C [48]; see also Fig.
+S9 for a simulated LT-TERS spectrum [46]).
+The slight discrepancy between the experimental and calculated frequencies are consistent with the typical spurious softening of the PBE functional.
+Compared to the calculated gas-phase value, the vibrational frequency shows a red-shift of 73 cm-1 upon the adsorption, which is larger than, but in reasonable agreement with the experiment.
 The TERS intensity is strongly correlated with the localized surface plasmonic resonance (LSPR) of the tip–sample junction.
-In general, there are two types of enhancement mechanisms effects in TERS [30, 38]: electromagnetic enhancement (originating from the LSPR of the junction with a plasmonic tip acting as an antenna for both incident light and Raman scattering) and chemical enhancement [originating from the interactions between the target molecule and the tip/surface].
+In general, two types of enhancement mechanisms contribute to TERS [36, 47]: electromagnetic enhancement through LSPR and chemical enhancement originating from charge transfer between the target molecule and the tip/surface.
 In physisorption systems, the electromagnetic enhancement is expected to be dominant.
-We examined the excitation wavelength dependence (Fig.
-S10), verifying the contribution of LSPR to the TERS intensity; when the scattering light wavelength matches with the resonance range of the LSPR, strong TERS peak was observed.
-The electromagnetic enhancement strongly depends on the cavity size.
-Figure 2(a) shows the gap-distance dependent TERS spectra recorded for D2/Ag(111).
-When the tip approaches the surface, the TERS intensities of the rotational and vibrational modes increase exponentially [Fig.
-2(b)].
+To elucidate the role of electromagnetic enhancement, it is crucial to investigate the gap-distance dependence of TERS signals.
+Figure 2(a) shows the waterfall plot of the gap-distance dependent TERS spectra recorded for D2/Ag(111).
+When the tip approaches the surface, the TERS intensities of the rotational and vibrational modes increase apparently exponentially [Fig.
+2(b)], in agreement with the prediction based on the electric field of a light-induced dipole in the plasmonic tip [57].
 A broad background with a maximum at 590 nm is attributed to electronic Raman scattering which represents the spectral response of the LSPR in the junction [58].
-The gap-distance dependent intensities of both rotational and vibrational modes are in good agreement with that of the plasmonic feature [Fig.
+The gap-distance dependent intensities of both rotational and vibrational modes closely follow the plasmonic feature [Fig.
 2(b)], indicating the dominant contribution of the gap-mode plasmon to the TERS enhancement.
+Additionally, we observed that changes in both the plasmon background and the molecular Raman peak occurred by an accidental modification of the tip-apex structure (Fig.
+S10), implying that the TERS enhancement is susceptible to the atomic-scale structure of the tip apex, which generates a plasmonic picocavity.
 The gap-distance dependence of the TERS intensities is well reproduced by simulating the electronic field distribution in the Ag-tip–Ag-surface junction using the finite element method (FEM) (see SM [48] for details).
-In the model, we assume a nanotip (0.5-nm radius of curvature) attached to a blunt tip body [51].
+To model the plasmonic picocavity, we assume an atomic-scale protrusion (0.5-nm radius of curvature) attached to a blunt tip body [40].
 The black curve in Fig.
-2(c) shows the field enhancement sampling at the position of a D2 molecule located under the Ag tip, indicating an exponential enhancement as the tip approaches.
+2(c) shows the field enhancement at the expected position of a D2 molecule located beneath the Ag tip, indicating a steep increase as the tip approaches.
 We confirmed the critical role of this sub-nanometric structure in TERS; as shown by the orange curve in Fig.
-2(c), the tip without the nanotip exhibits a more moderate increase of the field enhancement as the tip approaches, which does not fit the experimental data [Fig.
+2(c), the tip without the atomic-scale protrusion exhibits a more moderate increase of the field enhancement as the tip approaches, which does not fit the experimental data [Fig.
 2(b)].
@@
 2(c), inset], indicating much lower enhancement.
-Therefore, TERS will have “nearly” single-molecule sensitivity, even for the physisorption system.
-The rotational/vibrational spectroscopy is sensitive to the interaction between a target molecule and a local environment.
-We examine how the picocavity influences the rotational/vibrational state of H2 and D2.
-Figures 3(a)–3(d) show the gap-distance dependence of high-resolution TERS spectra focusing on the rotational/vibrational modes of H2/D2; notably, only the vibrational frequency of H2 exhibits a significant change [Fig.
-3(c)].
-Figure 3(e) shows the TERS spectra at three tip heights marked by the arrows in Fig.
-3(c), indicating the red-shift and peak broadening as the gap distance decreases.
-Peak-fitting analysis for H2 [red dotted curve in Fig.
-3(f)] indicates a corresponding peak shift of 34 cm-1 with a 200 pm displacement of the tip, while it is only 3 cm-1 for D2 (blue dotted curve).
-On the other hand, the rotational modes for both H2 and D2 [red and blue solid curves in Fig.
-3(f), respectively] are almost unchanged, in good agreement with the calculated interatomic distance independent of the tip height as described above.
-The significant (almost negligible) change of the vibrational mode of H2 (D2) can be explained by the anharmonicity of the potential energy surface (PES) of the molecule in the picocavity.
-The mode softening (red-shift and broadening) is associated with the PES of the molecule–surface distance, which is modified by the gap distance through the vdW interaction with the tip.
-Red-/Blue-shift of the intramolecular stretching mode through the vdW interaction has been reported by far-field Raman spectroscopy of hydrogen molecules physisorbed on carbon nanotubes [52].
-The PES modification by the tip proximity was already reported for Au-tip–H2–Au-surface junction [53]; however, we clarified that the intermolecular interactions also play an important role in the PES shape (Figs.
-S4–S6 [39]).
-Since H2 has a higher zero-point energy (ZPE) than D2, their vibrational frequencies can be more affected by the PES modification.
-Anharmonicity may also modify the lifetime of molecular vibrations [54, 55], which could lead to the peak broadening.
-However, the influence of neighboring molecules may not be negligible [Fig.
-2(c)], which would contribute to a shoulder at a lower frequency side of the vibrational peak [Fig.
-3(e)].
-To gain further insights into the observed red-shift, we calculated the PES experienced by the H2/D2 molecules directly under the tip at several gap distances .
-The PES is obtained by calculating the total energies by moving the molecule at several molecule–surface distances in the gap while the other atoms are fixed after the DFT optimization.
-The calculated PES exhibits a transition from a Morse-like potential [Fig.
-3(h)] to an asymmetric double-well potential [Fig.
-3(j)] as the gap distance decreases (see also Figs.
-S4 and S5).
-The quantum-corrected position for the H2/D2 molecule directly under the tip is obtained by solving the one-dimensional Schrödinger equation in the sinc-function discrete-variable representation (see detailed method, Fig.
-S6, and Table S1 in SM [39]).
-In parallel, we calculated the vibrational frequency of H2/D2 at each and with the i-PI code [56].
-We found a hallmark of vibrational coupling, i.e., -dependent vibrational frequencies (Figs.
-S7 and S8).
-Figure 3(g) [3(i)] shows at = 0.94 [0.74] nm, where the stretching frequency is calculated to be 4247 [4230] cm-1.
-At the far [close] tip distance, is calculated to be 320 [341] pm, while = 316 [333] pm, as shown in Fig.
-3(h) [3(i)].
-This indicates that D2 is closer to the surface and farther from the tip than H2, due to the ZPE difference, and that both H2 and D2 molecules are slightly attracted toward the tip by the change from 0.94 to 0.74 nm.
-By the tip proximity, we predict 18 cm-1 red-shift for H2, but only 9 cm-1 red-shift for D2.
-This result is qualitatively consistent with the experiment, but under(over)estimates the red-shift for H2 (D2).
-The absolute gap distance may be one of the possible reasons for this discrepancy; the gap distance in the experiment will be slightly different between H2 and D2 even with the same STM setpoint because of different tip-apex geometries.
-Finally, because the calculations regard the quantum-mechanical electronic ground-state only, we cannot rule out that non-adiabatic effects in the junction could further influence the observed isotopic differences in the TERS spectra.
-Such an analysis will be the subject of future work.
-With LT-TERS, we observed the rotational/vibrational transitions of single hydrogen molecules through the interaction with the picocavity field.
-Our results suggest that LT-TERS is applicable to weakly adsorbed molecules, offering deeper insights into the local structures, reactions, and dynamics of adsorbates, such as chemical reactions at active sites [57] and surface diffusion (spillover) in catalytic processes [58], at the single-molecule level.
+Therefore, TERS can achieve ‘nearly’ single-molecule sensitivity, even for physisorption systems.
+The contribution of LSPR to the TERS intensity was further confirmed by our excitation-wavelength dependent measurements (SM Sect.
+II-B [48]); a strong TERS peak was observed only when the scattering light wavelength matches with the resonance range of the LSPR (Fig.
+S3).
+TERS reveals the sensitivity of the rotational/vibrational states to the interactions between a target molecule and its local environment.
+Figures 3(a)–3(d) show the gap-distance dependence of high-resolution TERS spectra focusing on the rotational/vibrational modes of H2/D2.
+Notably, only the vibrational frequency of H2 exhibits a significant red-shift as the gap distance decreases [Fig.
+3(c)]; H2 shows a peak shift of 33 cm-1 with a 200 pm displacement of the tip, whereas is merely cm-1 for D2.
+Additionally, the peak width of H2 vibrational mode also increases as the gap distance decreases (see SM Sect.
+II-C [48] for details).
+In contrast, the peak positions of the rotational modes for both H2 [Fig.
+3(a)] and D2 [3(b)] are almost unchanged, consistent with our calculations showing that the interatomic distance is independent of the tip height, as described above.
+The isotope dependence of the red-shift of the vibrational band far exceeds the scale of typical isotope effects.
+To understand this notable behavior, we calculated the potential energy profile along the rigid vertical displacement of the center of mass of one molecule between the surface and the tip apex [Fig.
+4(a)], while varying the tip–sample distance [see profiles (i)–(iv) at different in Fig.
+4(b)].
+At each position, we calculated the H2 and D2 harmonic intramolecular vibrational frequency (see SM Sect.
+II-C [48]).
+The calculations indicate a pronounced coupling of the stretch coordinate with the molecular translation along , evidenced by a steep variation of along this coordinate (Fig.
+S8).
+Furthermore, this variation changes depending on the tip–sample distance, leading us to define .
+As detailed in SM Sect.
+I-C [48], for each , we constructed a model 2D Schrödinger equation in the coupled coordinates and , parameterized by and .
+The solution of this model gives us the expected intramolecular stretch vibrational transition energy for each isotope, fully accounting for the anharmonic coupling between these coordinates.
+We concluded that if both H2 and D2 probe the same well of [Fig.
+4(b)], the isotope effect on can account at most for a factor 2 between H2 and D2 (Tables S2 and S3), in disagreement with the experiment (Fig.
+3).
+We demonstrate that this discrepancy can be resolved by accounting for density differences between H2 and D2.
+At lower temperatures, the solids formed by H2 and D2 have different molar volumes due to nuclear quantum effects [zero-point energy (ZPE) swelling [59]; see also SM Sect.
+I-B [48]].
+As a result, solid D2 is more than twice as dense as solid H2 at 4.2 K [60].
+Because the interaction potential between the molecules and the surface exhibits almost no corrugation [52], we expect that H2 adsorbs on Ag(111) at 10 K more sparsely than D2.
+To explore a limiting low-density regime, we repeated the calculations discussed in the previous paragraph for a coverage of 0.06 ML of H2 [Fig.
+4(c)].
+As shown in Fig.
+4(d), the reduced intermolecular interactions at the surface cause the minimum of along to lie closer to the tip instead of closer to the surface, as decreases [(i) to (iv)].
+Consequently, the average position of the molecule center of mass [bullet points in Figs.
+4(b) and (d)] also shifts towards the tip as decreases [dotted line in Fig.
+4(d)].
+Computing in this scenario with the 2D quantum model of and , we find a larger variation with than in the higher-density case (Table S2).
+As shown in Fig.
+4(e), considering the higher-density scenario for D2 (the filled bullets) and the lower-density scenario for H2 (the empty bullets) can account for the large isotope effect in the red-shift of the stretch transition as the tip approaches.
+Figure 4(e) also shows the experimental values from the plots in Fig.
+3 (the solid and dotted curves for H2 and D2, respectively; see also Fig.
+S11).
+Although the exact gap distance in STM experiment is hard to determine (see SM Sect.
+II-C [48]), the overlap of the experimental curves with the calculated plots in Fig.
+4(e) shows good agreement.
+The remaining numerical discrepancies can be attributed to the limitation of the DFT level of theory (see SM Sect.
+I-B [48]), the exact tip apex geometry, and to the uncertainty in determining the exact local surface coverage.
+Nevertheless, our results highlight the crucial role of intermolecular interactions in accurately interpreting vibrational spectroscopy of hydrogen isotopes in nanojunctions, where nuclear quantum effects have a significant impact [13].
+In summary, with LT-TERS, we observed the rotational/vibrational transitions of local H2 and D2 molecules through the interaction with the picocavity field.
+Our results show that LT-TERS is applicable to weakly adsorbed molecules, offering deeper insights into the local structures, reactions, and dynamics of adsorbates, such as chemical reactions at active sites [61] and surface diffusion (spillover) in catalytic processes [62], at the single-molecule level.
 Acknowledgements.
@@
 - Yu et al. [1983] C.-f. Yu, K. B. Whaley, C. S. Hogg, and S. J. Sibener, Selective adsorption resonances in the scattering of -H2, -H2, -D2 and -D2 from Ag(111), Phys. Rev. Lett. 51, 2210 (1983).
+- Sakurai et al. [1988] M. Sakurai, T. Okano, and Y. Tuzi, Ortho-para conversion of -H2 physisorbed on Ag(111) near two-dimensional condensation conditions, Appl. Surf. Sci. 33, 245 (1988).
 - Anger et al. [1989] G. Anger, A. Winkler, and K. Rendulic, Adsorption and desorption kinetics in the systems H2/Cu(111), H2/Cu(110) and H2/Cu(100), Surf. Sci. 220, 1 (1989).
+- Williams et al.
+[2002] K.
+A.
+Williams, B.
+K.
+Pradhan, P.
+C.
+Eklund, M.
+K.
+Kostov, and M.
+W.
+Cole, Raman spectroscopic investigation of H2, HD, and D2 physisorption on ropes of single-walled, carbon nanotubes, Phys.
+Rev.
+Lett.
+88, 165502 (2002).
 - Fukutani et al.
@@
 - Niki et al. [2008] K. Niki, T. Kawauchi, M. Matsumoto, K. Fukutani, and T. Okano, Mechanism of the ortho-para conversion of hydrogen on Ag surfaces, Phys. Rev. B 77, 201404(R) (2008).
+- Sugimoto and Fukutani [2014] T. Sugimoto and K. Fukutani, Effects of rotational-symmetry breaking on physisorption of ortho-and para-H2 on Ag(111), Phys. Rev. Lett. 112, 146101 (2014).
+- Sugimoto et al. [2017] T. Sugimoto, Y. Kunisada, and K. Fukutani, Inelastic electron tunneling mediated by a molecular quantum rotator, Phys. Rev. B 96, 241409(R) (2017).
 - Gupta et al. [2005] J. Gupta, C. Lutz, A. Heinrich, and D. Eigler, Strongly coverage-dependent excitations of adsorbed molecular hydrogen, Phys. Rev. B 71, 115416 (2005).
@@
 B 97, 195417 (2018).
+- Merino et al.
+[2018] P.
+Merino, A.
+Rosławska, C.
+C.
+Leon, A.
+Grewal, C.
+Große, C.
+González, K.
+Kuhnke, and K.
+Kern, A single hydrogen molecule as an intensity chopper in an electrically driven plasmonic nanocavity, Nano Lett.
+19, 235 (2018).
 - Wang et al. [2022] L. Wang, Y. Xia, and W. Ho, Atomic-scale quantum sensing based on the ultrafast coherence of an H2 molecule in an STM cavity, Science 376, 401 (2022).
@@
 - Trouwborst et al. [2009] M. Trouwborst, E. Huisman, S. van der Molen, and B. van Wees, Bistable hysteresis and resistance switching in hydrogen-gold junctions, Phys. Rev. B 80, 081407(R) (2009).
+- Haran [2010] G. Haran, Single-molecule Raman spectroscopy: a probe of surface dynamics and plasmonic fields, Acc. Chem. Res. 43, 1135 (2010).
 - Ding et al.
@@
 Aizpurua, Atomic-scale lightning rod effect in plasmonic picocavities: A classical view to a quantum effect, ACS Nano 12, 585 (2018).
+- Liu et al.
+[2019] S.
+Liu, M.
+Müller, Y.
+Sun, I.
+Hamada, A.
+Hammud, M.
+Wolf, and T.
+Kumagai, Resolving the correlation between tip-enhanced resonance raman scattering and local electronic states with 1 nm resolution, Nano Lett.
+19, 5725 (2019).
+- Liu et al.
+[2023] S.
+Liu, F.
+P.
+Bonafe, H.
+Appel, A.
+Rubio, M.
+Wolf, and T.
+Kumagai, Inelastic light scattering in the vicinity of a single-atom quantum point contact in a plasmonic picocavity, ACS Nano 17, 10172 (2023).
+- Park et al.
+[2024] Y.
+Park, I.
+Hamada, A.
+Hammud, T.
+Kumagai, M.
+Wolf, and A.
+Shiotari, Atomic-precision control of plasmon-induced single-molecule switching in a metal–semiconductor nanojunction, Nat.
+Commun.
+15, 6709 (2024).
 - Zhang et al.
@@
 - Lee et al. [2019] J. Lee, K. T. Crampton, N. Tallarida, and V. A. Apkarian, Visualizing vibrational normal modes of a single molecule with atomically confined light, Nature 568, 78 (2019).
-- Liu et al.
-[2023] S.
-Liu, F.
-P.
-Bonafe, H.
-Appel, A.
-Rubio, M.
-Wolf, and T.
-Kumagai, Inelastic light scattering in the vicinity of a single-atom quantum point contact in a plasmonic picocavity, ACS Nano 17, 10172 (2023).
 - Xu et al.
@@
 Hou, Determining structural and chemical heterogeneities of surface species at the single-bond limit, Science 371, 818 (2021).
+- Litman et al.
+[2023] Y.
+Litman, F.
+P.
+Bonafé, A.
+Akkoush, H.
+Appel, and M.
+Rossi, First-principles simulations of tip enhanced Raman scattering reveal active role of substrate on high-resolution images, J.
+Phys.
+Chem.
+Lett.
+14, 6850 (2023).
 - Chen and Jensen [2023] R. Chen and L. Jensen, Interpreting chemical enhancements of surface-enhanced Raman scattering, Chem. Phys. Rev. 4 (2023).
@@
 [63, 40, 49, 50, 51, 64, 46, 41, 52, 65, 17, 66, 67, 54, 68, 57, 69, 4, 56, 8, 11, 14, 15, 9, 70, 71, 18].
-- Liu et al.
-[2020] S.
-Liu, B.
-Cirera, Y.
-Sun, I.
-Hamada, M.
-Müller, A.
-Hammud, M.
-Wolf, and T.
-Kumagai, Dramatic enhancement of tip-enhanced Raman scattering mediated by atomic point contact formation, Nano Lett.
-20, 5879 (2020).
 - Blum et al.
@@
 23, 7875 (2021).
-- Litman et al.
-[2023] Y.
-Litman, F.
-P.
-Bonafé, A.
-Akkoush, H.
-Appel, and M.
-Rossi, First-principles simulations of tip enhanced Raman scattering reveal active role of substrate on high-resolution images, J.
-Phys.
-Chem.
-Lette.
-14, 6850 (2023).
-- Pockrand and Otto [1981] I. Pockrand and A. Otto, Surface enhanced raman scattering (SERS): Annealing the silver substrate, Solid State Commun. 38, 1159 (1981).
+- Cirera et al.
+[2022] B.
+Cirera, Y.
+Litman, C.
+Lin, A.
+Akkoush, A.
+Hammud, M.
+Wolf, M.
+Rossi, and T.
+Kumagai, Charge transfer-mediated dramatic enhancement of Raman scattering upon molecular point contact formation, Nano Lett.
+22, 2170 (2022).
+- Cui et al.
+[2023] J.
+Cui, Y.
+Zhang, and Z.-C.
+Dong, Atomistic polarization model for Raman scattering simulations of large metal tips with atomic-scale protrusions at the tip apex, Nanophotonics 12, 4031 (2023).
 - Veirs and Rosenblatt [1987] D. K. Veirs and G. M. Rosenblatt, Raman line positions in molecular hydrogen: H2, HD, HT, D2, DT, and T2, J. Mol. Spectrosc. 121, 401 (1987).
-- Huber and Herzberg [1979] K. P. Huber and G. Herzberg, Molecular spectra and molecular structure: Constants of diatomic molecules (Van Nostrand, New York, 1979) pp. 716.
+- Pettinger et al.
+[2009] B.
+Pettinger, K.
+F.
+Domke, D.
+Zhang, G.
+Picardi, and R.
+Schuster, Tip-enhanced Raman scattering: influence of the tip-surface geometry on optical resonance and enhancement, Surf.
+Sci.
+603, 1335 (2009).
 - Kamimura et al. [2022] R. Kamimura, T. Kondo, K. Motobayashi, and K. Ikeda, Surface-enhanced electronic Raman scattering at various metal surfaces, Phys. Status Solidi B 259, 2100589 (2022).
-- Liu et al.
-[2019] S.
-Liu, M.
-Müller, Y.
-Sun, I.
-Hamada, A.
-Hammud, M.
-Wolf, and T.
-Kumagai, Resolving the correlation between tip-enhanced resonance raman scattering and local electronic states with 1 nm resolution, Nano Lett.
-19, 5725 (2019).
-- Williams et al.
-[2002] K.
-A.
-Williams, B.
-K.
-Pradhan, P.
-C.
-Eklund, M.
-K.
-Kostov, and M.
-W.
-Cole, Raman spectroscopic investigation of H2, HD, and D2 physisorption on ropes of single-walled, carbon nanotubes, Phys.
+- Markland et al. [2008] T. E. Markland, S. Habershon, and D. E. Manolopoulos, Quantum diffusion of hydrogen and muonium atoms in liquid water and hexagonal ice, J. Chem. Phys. 128, 194506 (2008).
+- Megaw and Simon [1936] H. D. Megaw and F. Simon, Density and compressibility of solid hydrogen and deuterium at 4.2 K, Nature 138, 244 (1936).
+- Cai et al. [2023] Z.-F. Cai, N. Kumar, and R. Zenobi, Probing on-surface chemistry at the nanoscale using tip-enhanced Raman spectroscopy, CCS Chem. 5, 55 (2023).
+- Yang et al. [2024] J.-L. Yang, Y.-W. Zhou, M.-Y. Jia, Q. Liu, and Z.-F. Cai, Insights into heterogeneous catalysis by tip-enhanced Raman spectroscopy, ChemCatChem 16, e202400370 (2024).
+- Summers [1969] R. L. Summers, Empirical observations on the sensitivity of hot cathode ionization type vacuum gages, NASA Technical Report NASA-TN-D-5285 (1969).
+- Hermann and Tkatchenko [2020] J.
+Hermann and A.
+Tkatchenko, Density functional model for van der Waals interactions: Unifying many-body atomic approaches with nonlocal functionals, Phys.
 Rev.
 Lett.
-88, 165502 (2002).
-- Sugimoto et al. [2017] T. Sugimoto, Y. Kunisada, and K. Fukutani, Inelastic electron tunneling mediated by a molecular quantum rotator, Phys. Rev. B 96, 241409(R) (2017).
-- Persson and Ryberg [1989] B. Persson and R. Ryberg, Vibrational line shapes of low-frequency adsorbate modes: CO on Pt(111), Phys. Rev. B 40, 10273 (1989).
-- Ueba [1986] H. Ueba, Vibrational lineshapes of adsorbates on solid surfaces, Prog. Surf. Sci. 22, 181 (1986).
+124, 146401 (2020).
 - Kapil et al.
@@
 236, 214 (2019).
-- Cai et al. [2023] Z.-F. Cai, N. Kumar, and R. Zenobi, Probing on-surface chemistry at the nanoscale using tip-enhanced Raman spectroscopy, CCS Chem. 5, 55 (2023).
-- Yang et al. [2024] J.-L. Yang, Y.-W. Zhou, M.-Y. Jia, Q. Liu, and Z.-F. Cai, Insights into heterogeneous catalysis by tip-enhanced Raman spectroscopy, ChemCatChem 16, e202400370 (2024).
-- Summers [1969] R. L. Summers, Empirical observations on the sensitivity of hot cathode ionization type vacuum gages, NASA Technical Report NASA-TN-D-5285 (1969).
-- Hermann and Tkatchenko [2020] J.
-Hermann and A.
-Tkatchenko, Density functional model for van der Waals interactions: Unifying many-body atomic approaches with nonlocal functionals, Phys.
-Rev.
-Lett.
-124, 146401 (2020).
 - Schwartz [1985] C. Schwartz, High-accuracy approximation techniques for analytic functions, J. Math. Phys. 26, 411 (1985).
@@
 - Jiang et al. [2016] Y. Jiang, S. Pillai, and M. A. Green, Realistic silver optical constants for plasmonics, Sci. Rep. 6, 30605 (2016).
-- Pettinger [2010] B. Pettinger, Single-molecule surface-and tip-enhanced Raman spectroscopy, Mol. Phys. 108, 2039 (2010).
 - Wolniewicz [1983] L. Wolniewicz, The state vibration-rotational energies of the H2, HD, and D2 molecules, J. Chem. Phys. 78, 6173 (1983).
-- Sakurai et al. [1988] M. Sakurai, T. Okano, and Y. Tuzi, Ortho-para conversion of -H2 physisorbed on Ag(111) near two-dimensional condensation conditions, Appl. Surf. Sci. 33, 245 (1988).
 - Teal and MacWood [1935] G. K. Teal and G. E. MacWood, The Raman spectra of the isotopic molecules H2, HD, and D2, J. Chem. Phys. 3, 760 (1935).
+- Persson and Ryberg [1989] B. Persson and R. Ryberg, Vibrational line shapes of low-frequency adsorbate modes: CO on Pt(111), Phys. Rev. B 40, 10273 (1989).
 Supplemental Material: Picocavity-enhanced Raman Spectroscopy of Physisorbed H2 and D2 Molecules
@@
 The experiments were performed in an ultrahigh vacuum (UHV) chamber equipped with an STM-based TERS machine (modified UNISOKU USM-1400) at a sample temperature of 10 K.
-A single-crystalline Ag(111) surface (MaTeck) was cleaned by Ar-sputtering-and-annealing cleaning cycles in the chamber.
-H2 (99.999%, Westfalen AG) or D2 (99.7%, Air Liquide) gas was dosed into the STM unit cooled at 10 K via a variable leak valve.
-The gas exposure (Langmuir; L) is estimated by , where [ Torr], [s], and denote the displayed pressure value of the gauge (corrected by N2), the exposure duration, and the correction factor for the gauge.
-A Bayard-Alpert ionization gauge is located outside the cryogen-cooled STM unit in the chamber and (0.388) is used for H2 (D2) [59].
-We used a chemically-etched Ag tip that was sharpened by focused-ion-beam milling as described elsewhere [40].
+A single-crystalline Ag(111) surface (MaTeck) was cleaned by Ar-sputtering-and-annealing cycles in the chamber.
+H2 (99.999%, Westfalen AG) or D2 (99.7%, Air Liquide) gas was introduced into the STM unit cooled at 10 K via a variable leak valve.
+The gas exposure (Langmuir; L) is estimated by , where [ Torr], [s], and denote the displayed pressure value of the gauge (corrected by N2), the exposure duration, and the correction factor for the gauge, respectively.
+A Bayard-Alpert ionization gauge is located outside the cryogen-cooled STM unit in the chamber and we used for H2 and for D2 [63].
+We used a chemically-etched Ag tip that was further sharpened by focused-ion-beam milling as described elsewhere [40].
 The bias voltage was applied to the sample while the tip was grounded.
 The STM images were obtained in the constant current mode at = 50 pA.
-The cleanness of the cleaned surface and tip was confirmed by STM images [Fig.
+The cleanness of the Ag(111) surface and tip was confirmed by STM images [Fig.
 S1(a)].
 We also confirmed that a TERS spectrum of the Ag–Ag junction before hydrogen dosing had no molecule-derived peaks (Fig.
-S2, at 0 L)
-In the STM junction, -polarised visible laser (532-nm solid-state or 633-nm HeNe) was incident through a fused silica window and a Ag-coated parabolic mirror (numerical aperture of 0.6).
-The incident laser power is 5.9 (9.7) mW at = 532 (633) nm.
+S2, at 0 L).
+In the STM junction, -polarised visible laser (532-nm solid-state or 633-nm HeNe) was directed through a fused silica window and an Ag-coated parabolic mirror (numerical aperture of 0.6).
+The incident laser power is 5.9 mW at = 532 nm and 9.7 mW at 633 nm.
 The parabolic mirror is mounted on a five-axis piezoelectric-motor stage in the LT-STM unit and the laser is focused at the tip apex by adjusting the position and angle of the parabolic mirror.
-The scattered light corrected by the parabolic mirror was directed to a spectrometer (AndorShamrock 303i) outside the UHV chamber via a beam splitter and a longpass filter specific for each excitation wavelength.
+The scattered light collected by the parabolic mirror was directed to a spectrometer (AndorShamrock 303i) outside the UHV chamber via a beam splitter and a longpass filter, each selected according to the excitation wavelength.
 For the TERS and STM-induced luminescence (STML) spectra, we used the gratings 150 (for Fig.
@@
 1, 2, and S2), and 1200 line/mm (Fig.
-3).
-A sparser grating allows the acquisition of wider-wavelength range spectra, as shown in Fig.
-S3 (see Sec.
-II-B), but Raman peaks become broad and asymmetric.
-Thus, for the high-frequency-resolution measurements shown in Fig.
+3 and S10).
+A lower-density grating allows the acquisition of spectra over a wider wavelength range, as shown in Fig.
+S3 (see Sect.
+II-B), but resulting in broad and asymmetric Raman peaks.
+Therefore, for the high-frequency-resolution measurements shown in Fig.
 3, the densest grating was used.
@@
 [51] in all simulations reported in the main text.
-Figure S4 shows a comparison of binding energies with the many-body NL-MBD dispersion correction of Ref.
+Figure S4 shows a comparison of binding energies with the many-body MBD-NL dispersion correction of Ref.
 [64], indicating a weaker binding energy by only 5 meV and a binding distance 0.1 Å closer to the surface.
@@
 All the other atoms were kept fixed at their bulk (or tip) positions.
-When we performed simulations including a model tip structure, this was modeled by a Ag19 cluster at varying distances from the surface.
-Although this tip size would not be sufficient to capture a proper plasmonic response of the tip, the tip structure can capture the local interactions close to the tip apex [46, 36].
+When we performed simulations including a model tip structure, this was modeled by an Ag19 cluster at varying distances from the surface.
+The geometry of the tip is shown in Fig.
+S5.
+Although this tip size would not be sufficient to capture a proper plasmonic response of the tip, the tip structure can capture the local interactions with molecules close to the tip apex [46, 54, 41].
 We inspected binding sites and binding energies of a single adsorbed H2 molecule, finding all values in good agreement with previous studies [52].
@@
 The model we used in most simulations contained 11 molecules in the unit cell (see rationale below).
-For this molecular concentration, we evaluated the intermolecular interaction by comparing the binding curves (orange and green curves in Fig.
+For this molecular concentration, we evaluated the intermolecular interaction by comparing the binding curves (orange curve in Fig.
 S4) with that of the single-molecule case (blue); the molecule–molecule interaction in our models amounts to 55 meV/molecule.
@@
 We considered a full monolayer (1 ML), where all fcc-hollow sites of the unit cell were occupied by H2.
-With this setup we performed Born-Oppenheimer molecular dynamics (BOMD) and path-integral molecular dynamics (PIMD) at LT ( = 20 K) with the i-PI code [56] connected to FHI-aims.
+With this setup we performed Born-Oppenheimer molecular dynamics (BOMD) and PIMD at LT ( = 20 K) with the i-PI code [65] connected to FHI-aims.
 PIMD simulations used 12 beads in conjunction with a normal-mode colored-noise thermostat with 8 auxiliary variables and .
@@
 While the full monolayer was stable during more than 10 ps on the classical trajectories, in the quantum trajectory 5 molecules readily desorbed (within the first picosecond) from the substrate.
-For the quantum case we validated this observation by simulating trajectories with a parameterized Lennard-Jones potential representing the Ag(111) surface.
+For the quantum case, we validated this observation by simulating trajectories with a parameterized Lennard-Jones potential representing the Ag(111) surface.
 We could calculate the radius of gyration of H atoms from the PIMD simulations, belonging to the 3 molecules closest to the tip at any point in the simulation.
 For the closest tip distance considered ( = 7.4 Å), these were 0.33 Å along and and 0.30 Å along ( = 20 K), indicating a small anisotropy.
-Converting these numbers to an effective thermal de Broglie wavelength and applying a inverse-square-root scaling for the temperature and mass, we would obtain Å and Å at 10 K.
-We note that, depending on the density of H2 molecules on the surface, is about the length of the inter-particle spacing at 10 K and signatures of bosonic exchange could begin to be observed.
+Converting these numbers to an effective thermal de Broglie wavelength and applying a inverse-square-root scaling for the temperature and mass, we would obtain Å for H2 and Å for D2 at 10 K.
+We note that, depending on the density of H2 molecules on the surface, for H2 is about the length of the inter-particle spacing at 10 K and signatures of bosonic exchange could begin to be observed.
 These features are disregarded in the current simulations.
-Figure S5 shows the energy profiles of a H2 molecule with varying distance to the Ag(111) surface at different tip-surface distances.
-We only varied the position of the molecule closest to the tip apex, while the others were kept at equilibrium [see also Figs.
-3(g) and 3(i) in the main text].
-The profile shapes are similar to an asymmetric double-well potential at closer tip–surface distances, while they assume a shape closer to a Morse potential at larger distances, which asymptotically approaches the profile of a H2 molecule on Ag(111) without the Ag tip (yellow curve in Fig.
-S5).
+Further, we note that even though these simulations provide a clear indication that quantum mechanical effects affect the density of the molecular coverage of the surface, a complete determination would require much longer simulations, which would take years to complete or require the development of a tailored machine-learning potential that surpasses the scope of this work.
+Figure S6 shows the energy profiles experienced by an H2 molecule when varying the distance of its center of mass with respect to the Ag(111) surface, at different tip-surface distances, and at 0.69 ML coverage.
+We only varied the position of the molecule closest to the tip apex, while the others were kept at equilibrium.
+The profile shapes are similar to an asymmetric double-well potential at closer tip–surface distances, while they assume a shape closer to a Morse potential at larger distances, which asymptotically approaches the profile of an H2 molecule on Ag(111) without the Ag tip (yellow curve in Fig.
+S6).
 The asymmetry of the double-well is strongly impacted by the presence of other molecules on the surface (molecule–molecule interactions).
-Figure S6 shows a comparison with the limiting case of a single molecule in the unit cell at a tip–surface distance of 9.4 Å.
+Figure S7(a) shows a comparison with the limiting case of a single molecule in the unit cell (0.06 ML coverage) at a tip–surface distance of 9.4 Å.
 In that case (orange curve), if the molecule can overcome the barrier, it will have a higher probability of being found closer to the tip, which was predicted in a Au-tip–single-H2–Au-surface junction [17].
-I.3 C. Exact DVR Models and Vibrational Analysis
-For the potentials shown in Fig.
-S5, we solved the one-dimensional Schrödinger equation exactly in the sinc-function discrete variable representation (DVR) [61, 62].
-We set the mass of the hydrogen atom to and that of deuterium to .
-In all cases, calculations were performed on a regularly spaced grid of 100 points, spanning values of from to with a potential-energy cut-off .
-The potential energies used in the DVR were obtained from a cubic spline interpolation of the DFT-calculated values in Sec.
-I-B.
-The resulting ZPEs and average ground-state displacements of the molecular center of mass from equilibrium,
-are given in Table S1, as well as those for D2.
-Here, is the molecular height considering the quantum mechanical ZPE contributions at the PES (i.e., the “quantum-corrected” molecular heights), and are the underlying ground-state wavefunctions, which are plotted in Fig.
-S7 to further illustrate the influence of tip–surface separation on the ZPEs and the equilibrium displacements.
-no tip | ||||||
-In parallel we conducted harmonic vibrational analysis with the i-PI code, varying the height of the H2/D2 molecule directly under the tip apex and varying the tip–sample distance [see also Figs.
-3(g) and 3(i)].
-Figures S9 and S9 show the dependence of the H–H and D–D stretching vibration on the mode corresponding to the vertical hindered translation of the adsorbed molecule, i.e., H2–Ag(111) and D2–Ag(111), respectively.
-At each , there is a red-shift of the interatomic stretching frequency with approaching tip, which follows a mass scaling between H2 and D2.
-As described above, with the DVR solutions, we obtain the the quantum-corrected molecule–surface distance .
-Based on the -dependent vibrational frequency curves, we obtain the vibrational frequency at , as shown by the blue (green) squares for H2 (D2) in Fig.
-S9 (S9).
-The plot indicates an anomalous isotopic red-shift for H2 (factor 2, instead of ), stemming from the two following factors: (1) strong anharmonic character of the hindered translation of the molecule, which is dominated by the vdW interactions with the tip and the surface (see Fig.
-S5), and (2) the coupling of this mode with the vibrational stretching mode.
-We stress that different heights of the tip and different concentrations of the molecules can exacerbate this effect.
-The former because the red-shift is more pronounced at closer tip distances than at larger tip distances.
-The latter because weaker binding energies to the surface at lower concentrations result in softer potentials that assume a double-well character at larger tip distances (see the orange curve in Fig.
-S6).
-This effect can result in a more pronounced negative slope of the dashed curves in Figs.
-S9 and S9.
-It is challenging to determine the local molecular concentration closer to the tip apex at any measurement and a deeper study of of the vibrational frequency variation could provide indirect hints to the observed concentration.
-Although we neglected the anharmonic character of the H–H stretching coordinate in this study, we expect this effect not to change the qualitative discussion and introduce only rigid red-shifts on the values reported here.
+As we discuss in the main text, this effect is essential to explain the isotope-dependent red-shift observed experimentally.
+Figure S7(b) shows a comparison of the 0.69 ML coverage with the tip fixed at Å and different vdW dispersion corrections, added to the PBE functional.
+It is clear that details of the vdW dispersion impact the shape of the potential.
+In this case, the many-body vdW dispersion decreases the distance between the two minima of the asymmetric double-well
+I.3 C. Model 2D Schrödinger Equation
+[…]
+where is the potential energy of a molecule with its center of mass fixed at a distance above the substrate surface and with all other coordinates kept at their equilibrium values, as described at the end of Sec.
+I.2.
+The vibrational motion is parametrized by the reduced mass and angular frequency .
+We set the mass of a hydrogen atom to and use .
+The vibrational frequencies are computed from numerical (finite-difference) second derivatives of the DFT energies with respect to the vibrational coordinate,
+[…]
+We computed the numerically exact solutions to the 2D Schrödinger equation corresponding to Eq.
+(S1) using the sinc-function discrete variable representation (DVR) [66, 67].
+The equation was solved for high coverage (0.69 ML) with and .
+Given the large separation in the timescales of translational and vibrational motion, we also considered the vibrationally adiabatic approximation, whereby one fixes the translational coordinate in Eq.
+(S1) and solves the corresponding one-dimensional (1D) Schrödinger equation in , yielding the analytical harmonic vibrational energies
+These are then used to construct the vibrationally adiabatic translational potential,
+for which we solve the 1D Schrödinger equation numerically using DVR.
+The energy difference between the ground state solutions for and is an excellent approximation to the vibrational excitation energy of the 2D model (see Table S1).
+Given the accuracy of the vibrationally adiabatic approximation, we adopt it in all subsequent calculations.
+2D | ||||
+1D |
+[…]
+11.4 | ||||||
+9.4 | ||||||
+7.9 | ||||||
+7.4 |
+11.4 | ||||||
+9.4 | ||||||
+7.9 | ||||||
+7.4 |
 I.4 D. TERS Spectrum Simulation
-We conducted TERS simulations according to the methodology proposed in Ref.
+We conducted TERS simulations according to the methodology proposed in Refs.
 [46].
-We used Tip A discussed in the reference and calculated the anharmonic Raman spectrum based on a classical ab initio molecular dynamics (AIMD) trajectory (5 ps) of a molecular layer adsorbed on the surface and a tip apex located 4 Å above the molecular layer at = 20 K.
+We used Tip A discussed in Ref.
+[54] and calculated the anharmonic Raman spectrum based on an ab initio molecular dynamics (AIMD) trajectory (5 ps) of a molecular layer adsorbed on the surface and a tip apex located 4 Å above the molecular layer at = 20 K.
 We calculated the TERS intensity as , where was obtained by a local polarizability calculation at each geometry in the time-series given by the molecular dynamics.
-For this part of the calculation we only considered the free-standing molecular layer (no surface) at the positions adopted in the AIMD trajectory including the surface.
+For this part of the calculation, we only considered the free-standing molecular layer (no surface) at the positions adopted in the AIMD trajectory including the surface.
 We note that the method of Ref.
-[46] is unable to capture the electromagnetic plasmonic enhancement in its totality.
-We do not observe any noteworthy features in the spectrum and it overall agrees with the experimental spectrum.
-The red-shift of the rotational band in the simulation is explained by the lack of consideration of the nuclear spin, making calculations reproduce only classical rotations that are red-shifted if compared to the quantized transition of para-H2.
-This simulation had only an exploratory character and achieving statistical convergence through the use of parametrized machine-learned potentials, plus the inclusion of accurate plasmonic enhancements is beyond the scope of this paper.
+[46] is unable to capture the plasmonic enhancement in its totality.
+We do not observe any noteworthy features in the calculated spectrum and it overall agrees with the experimental spectrum.
+The red-shift of the rotational band in the simulation is explained by the absence of nuclear spin in the simulations, making calculations reproduce only classical rotations that are red-shifted compared to the quantized transition of para-H2.
+This simulation is of exploratory character, as achieving statistical convergence through the use of parametrized machine-learned potentials along with accurate plasmonic enhancements is beyond the scope of this paper.
 I.5 E. FEM Simulations
-To evaluate the electric field in the STM junction, we performed finite element method (FEM) simulations using COMSOL Multiphysics (version 6.2 with the Wave Optics Module).
-We used the three-dimentional model with a Ag tip and Ag sample plate described in Supporting Information of Ref.
+To evaluate the electric field in the STM junction, we performed FEM simulations using COMSOL Multiphysics (version 6.2 with the Wave Optics Module).
+We used the three-dimentional model with an Ag tip and Ag sample plate described in Supporting Information of Ref.
 [40].
-As needed, at the apex of the Ag tip (a length of 300 nm, an effective radius of 30 nm), a half-sphere protrusion with a radius of 0.5 nm is attached as a “nanotip.” A -polarized plane wave with of 532 nm and an electric field of 1 V/m is incident with an angle of 35∘ relative to the sample surface.
+To reproduce a plasmonic picocavity in the STM junction, at the apex of the Ag tip (a length of 300 nm, an effective radius of 30 nm), an Ag half sphere with a radius of 0.5 nm is attached as an atomic-scale protrusion.
+For comparison, we also constructed another Ag tip model with an effective radius of 30 nm without the atomic-scale protrusion [see the orange dotted curve in Fig.
+2(c) in the main text].
+A -polarized plane wave with of 532 nm and an electric field of 1 V/m is incident with an angle of 55∘ relative to the sample normal.
 The perfectly matched layers surrounding the volume are set to absorb all outgoing waves.
@@
 The surface normal component of the electric field is sampled.
-A Ag flat plate is placed beneath the tip apex with a gap distance .
+An Ag flat plate is placed beneath the tip apex with a gap distance .
 Here we define the origin of the gap distance ( = 0) as the tip height when the Ag-tip apex forms point contact with the Ag surface.
-As described above, the quantum-corrected molecular height is calculated to be approximately 0.32–0.33 nm for D2/Ag(111).
+As shown in Table S3, the molecular height is calculated to be 0.32–0.33 nm for D2/Ag(111).
 In the definition, taking into account the atomic radius of Ag, the position of the molecule corresponds to 0.2 nm above the Ag surface plane in the FEM model.
-Although the calculated depends on the tip height [Figs.
-3(h) and 3(j) in the main text], in the FEM model, we used the constant height for the sampling point at any gap distances.
+Although the calculated depends on the tip height (Table S3), in the FEM model, we used the constant height for the sampling point at any gap distances.
 We confirmed that differences in the definition or the molecular position on the order of 0.1 nm do not result in qualitative differences in the field enhancement factor curves.
@@
 2c of the main text represent the plots sampling at (, , ) = (0, 0, 0.20) nm and (0.29, 0, 0.20) nm, respectively.
-The tip heights in the DFT calculations are comparable with the gap distances with the FEM simulations that reproduced the Raman enhancement [Fig.
-2(c) in the main text; the gap distances of 550 to 850 pm in the simulation correspond to the tip height of 0.8 to 1.1 nm, considering the Ag-atom radius].
+As shown in Fig.
+2, the gap distances with the FEM simulations pm to 850 pm well reproduced the Raman enhancement at the experimental tip height to 0 pm.
+Considering the Ag-atom radius, for the simulations (Fig.
+4 in the main text) is comparable to nm.
+Therefore, pm corresponds to nm, which was used for the alignment of the experimental and theoretical plots displayed in Fig.
+4(e) of the main text.
 II II. Supplemental Results and Discussion
@@
 para-H2 (nuclear spin quantum number = 0) is more stable than ortho-H2 ( = 1) by 14 meV [69].
-para- (ortho-) H2 has only even (odd) rotational quantum numbers .
+para-H2 has only even rotational quantum numbers , whereas and ortho-H2 has only odd .
 Although ortho-H2 is the predominant species in the gas phase at room temperature (a temperature equilibrium composition of 75%), the molecular adsorption on metals enhances the ortho-to-para conversion, leading to the dominance of the para species on the surfaces [4].
@@
 Only -D2 was detected in our TERS spectra, indicating again that para-D2 was converted on the surface.
-Although the dominance of para-H2 in the TERS spectra is consistent with a previous HREELS study of H2 on a Ag film (0.5–10 L) [8], some previous reports detect ortho-H2 on Ag(111).
-A HREELS study on H2/Ag(111) [66] shows that rotational mode peaks of both para-H2 and ortho-H2 were detected under a higher H2 pressure ( Torr) while para-H2 was dominantly detected under a lower pressure ( Torr).
-Photostimulated desorption studies (30 L) [12, 13] shows that ortho-H2 and para-D2 are alive on Ag(111) on the order of hundred seconds and that the conversion of D2 is slower than H2.
-To compare with the above-mentioned studies, we performed TERS measurements during exposing D2 to a clean Ag(111) surfaces, as shown in Fig.
+Although we introduced normal hydrogen gases (n-H2 and n-D2) into the chamber, we observed rotational/vibrational transitions only for para-H2 and ortho-D2 in the TERS spectra (Fig.
+1 in the main text).
+The dominance of para-H2 on Ag(111) in the TERS spectra ( = 10 K) is consistent with a previous HREELS study of H2 on an Ag film (exposures of 0.5–10 L, = 10 K) [8], whereas some previous reports detect ortho-H2 on Ag(111).
+A HREELS study on H2/Ag(111) at 6 K [11] shows that rotational mode peaks of both para-H2 and ortho-H2 were detected under a higher H2 pressure ( Torr) while para-H2 was dominantly detected under a lower pressure ( Torr).
+Photostimulated desorption studies [14, 15] shows that at an exposure of 30 L and a temperature of 6 K, ortho-H2 and para-D2 are alive on Ag(111) on the order of hundred seconds and that the conversion of D2 is slower than H2.
+To compare with the above-mentioned studies, we performed TERS measurements while exposing D2 to a clean Ag(111) surfaces, as shown in Fig.
 S2.
@@
 It should be noted that even if the gas pressure and exposure values are comparable, the molecular coverage varies depending on the chamber configuration.
-Nevertheless, we confirmed that ortho-H2 and para-D2 were not detected at any pressures and exposures we used (0.8–2 Torr; 102–103 Langmuir).
-The HREELS observation of the metastable isomer only at high exposures [66] may be attributed to be hydrogen multilayers, where the second layer does not interact with the Ag surface.
-In contrast, the narrow TERS gap for the LT-TERS measurement would suppress the formation of the multiple layers.
+Nevertheless, we confirmed that ortho-H2 and para-D2 were not detected at any pressures and exposures we used (0.8–2 Torr; 102–103 L).
+The HREELS observation of the metastable isomer at 6 K only at high exposures [11] may be attributed to be hydrogen multilayers, where the second layer does not interact with the Ag surface.
+For our LT-TERS measurement, in contrast, only a monolayer would be formed at the higher temperature, 10 K.
 Another possibility that the metastable isomer were not detected in TERS is that the LSPR in the gap is accelerating the para-to-ortho conversion of the molecule under the tip.
 Although far-field 532-nm irradiation did not affect the conversion rate [15], the strong electric field localized in the STM junction could interact with the nuclear spin of the adsorbate.
-This effect would be verified through theoretical calculations.
 System | On Ag(111) | On Ag film | On Cu(001) | Gas phase | ||
@@
 To verify the contribution of LSPR to the TERS intensity, we examined the excitation wavelength () dependence.
-Figure S3(a) shows a STML spectrum recorded over a clean Ag(111) surface with a Ag tip.
+Figure S3(a) shows a STML spectrum recorded over a clean Ag(111) surface with an Ag tip.
 The LSPR results in a broad spectral feature in the visible range with a maximum at 560 nm.
@@
 The peak intensity of each mode significantly changes with , whereas the Raman shift value is unchanged.
-The vibrational intensity is weakened with = 633 nm, indicating the corresponding scattering light is not efficiently enhanced by the LSPR because the peak position ( = 779 nm) is located in the weak resonance range of the LSPR [Fig.
+The vibrational intensity is weakened with = 633 nm, indicating the corresponding Raman scattering is not efficiently enhanced by the LSPR because the peak position ( = 779 nm) is located in the weak resonance range of the LSPR [Fig.
 S3(a)].
 In contrast, both the rotational and vibrational peaks become intense with = 532 nm because they are more resonant with the LSPR.
+II.3 C. Tip-Height Dependence of Peak Intensity and Width
+To obtain the waterfall plots shown in Fig.
+3 of the main text, we conducted tip approach-and-retraction procedures while recording the TERS spectra [54, 41].
+Figure S10(a) is the full dataset for the H2 “vib” mode shown in Fig.
+3(b) of the main text.
+During recording TERS spectra, the tip height was lowered stepwise and then the tip was retracted from the surface.
+The symmetrical waterfall plot with respect to the closest tip height indicates the non-destructive measurement in the tip approach-and-retraction process.
+Figure S10(b) shows the horizontal line profiles of the waterfall plot for the H2 “vib” mode at the points indicated by the arrows in the same color.
+The three spectra indicates the red-shift and broadening of the H2 “vib” peak as the gap distance decreases.
+During recording the waterfall TERS plot, we occasionally observed a sudden intensity change caused presumably by the modification of the tip apex structure due to the strong tip–surface interaction at the small gap.
+Figure S10(c) shows an example, where a sudden drop of the TERS intensity was observed.
+The modification of the plasmon background intensity implies that the atomic-scale structure was changed by the strong attractive force exerted when the tip was brought close to the surface.
+At the same time as the plasmon feature, the D2 “rot” peak was also dropped, while the frequency of the “rot” mode is unchanged.
+This observation is in good agreement with the enhancement mechanism described in the main text (Fig.
+2); the picocavity-derived electric field plays a crucial role in the TERS intensity of physisorbed hydrogen molecules.
+To determine the experimental peak position shift as a function of the tip height in Fig.
+4(e) of the main text, the line profile of the waterfall plots at each tip height [Fig.
+3 in the main text; see also Fig.
+S10(b)] was fitted by a Gaussian peak with a constant background.
+The top and bottom panels of Fig.
+S11 shows the peak position and width, respectively, as a function of the tip height for the rotational (a) and vibrational (b) modes.
+The error bars indicate the standard deviation of the fitting.
+As the tip approaches, the peak intensity increases, reducing the error.
+The point at which the error becomes below 2 cm-1 by the tip approach was used as the reference point [the arrows in the top panel of Fig.
+S11(a) and (b)] and the peak position at this reference point was set to the origin of the peak shift (0 cm-1 of the top panel).
+For the H2 and D2 “vib” modes [top panel of Fig.
+S11(b)], the plots after the reference points were used as the experimental plots shown in Fig.
+4(e) of the main text.
+In Fig.
+4(e) of the main text, the experimental tip height pm is aligned to the tip–sample distance nm for the calculations.
+Although the accurate determination of the experimental tip–sample gap distance is difficult, this alignment is reasonable; as described in Sect.
+I-E, pm is comparable with pm and nm.
+As shown in Fig.
+S11(a), the frequency and width of the H2 and D2 “rot” peaks are almost unchanged by the tip.
+In contrast, the bottom panel Fig.
+S11(b) shows the strong peak broadening for the H2 “vib” mode [see also Fig.
+S10(b)] whereas the peak width for D2 “vib” is almost unchanged.
+A possible origin of the peak broadening is the anharmonicity, which can modify the lifetime of molecular vibrations [71].
+It is also entirely possible that if neighboring molecules influence the experimental signal [see Fig.
+2(c) in the main text], they would lead to the high-frequency shoulder of the vibration transition for H2, as these molecules would still lie on the surface [Fig.
+3(c)] and would therefore present a higher than the ones on the tip, which instead provide the most enhanced signal; however, this effect may not be pronounced because the coverage of is less than that of D2 (see Sect.
+I-C).
+Another possible origin if the movement of the molecule in the junction; because the double-well potential barrier is very small [see Fig.
+4(d) in the main text], the H2 molecule near the tip apex can move to a metastable site near the surface or adjacent adsorption sites on the surface [18] by a small voltage bias or plasmonic field applied in the STM junction.
+This movement instantaneously occurring during the acquisition time of the Raman spectrum can also modify the peak shape of the vibrational mode.
+We expect that the coexistence of these factors leads to the characteristic peak width depending on the tip height.

### v3

@@
 Abstract
-We report on tip-enhanced Raman scattering (TERS) of H2 and D2 molecules physisorbed within a plasmonic picocavity at 10 K.
-The intense Raman peaks resulting from the rotational and vibrational transitions are observed at sub-nanometer gap distances of the junction formed by an Ag tip and Ag(111) surface, where a picocavity-enhanced field plays a crucial role.
-A significant red-shift of the H–H stretch frequency is found as the gap distance decreases, while the D–D stretch frequency is unaffected.
+We report on tip-enhanced Raman spectroscopy of H2 and D2 molecules physisorbed within a plasmonic picocavity at 10 K.
+The intense Raman peaks resulting from the rotational and vibrational transitions are observed at subnanometer gap distances of the junction formed by an Ag tip and an Ag(111) surface, where a picocavity-enhanced field plays a crucial role.
+A significant redshift of the H-H stretch frequency is observed as the gap distance decreases, while the D-D stretch frequency is unaffected.
 Density functional theory, path-integral molecular dynamics, and quantum anharmonic vibrational energy calculations suggest that this unexpected isotope effect is explained by a different molecular density between H2 and D2 on the surface.
@@
 To understand and control the elementary processes in such physical and chemical phenomena, hydrogen adsorption has been intensively studied in surface science [5, 6, 7].
-Specifically, hydrogen, the lightest molecule, exhibits unique quantum effects that differ from those of other molecules when it is physisorbed on a surface through van der Waals (vdW) interactions [8, 9, 10, 11, 12, 13, 14, 15, 16, 17].
+Specifically, hydrogen, the lightest molecule, exhibits unique quantum effects that differ from those of other molecules when it is physisorbed on a surface through van der Waals interactions [8, 9, 10, 11, 12, 13, 14, 15, 16, 17].
 However, only a few methods are available for characterizing such physisorption systems as the weak adsorption energy ( meV) necessitates measurements at cryogenic temperatures.
-Low-temperature scanning tunneling microscopy (LT-STM) has been employed as a local characterization technique to investigate the dynamics of hydrogen molecules weakly adsorbed on surfaces [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28].
-Several studies have observed characteristic features in conductance spectra, attributed to the rotational transition of molecular hydrogen in the STM junction [21, 22, 23, 24].
+Low-temperature scanning tunneling microscopy has been employed as a local characterization technique to investigate the dynamics of hydrogen molecules weakly adsorbed on surfaces [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28].
+Several studies have observed characteristic features in conductance spectra, attributed to the rotational transition of molecular hydrogen in the scanning tunneling microscopy (STM) junction [21, 22, 23, 24].
 However, the interpretation remains controversial; the features within the corresponding energy range [30–60 meV (240–480 cm-1) for H2] could be attributed to configuration switching [18, 29, 30], to a molecule–substrate stretching mode [31], or to multiple phonon excitation [32].
@@
 Raman spectroscopy is a promising technique for characterizing hydrogen molecules on surfaces, as it allows the observation of both rotational and vibrational transitions of homonuclear diatomic molecules.
-Although the detection demands exceptional sensitivity due to the intrinsically small cross-section, single-molecule detection can be achieved by tip-enhanced and surface-enhanced Raman spectroscopy (TERS/SERS) [33, 34, 35, 36, 37].
-Recent studies have established that picocavities, atomic-scale protrusions present in plasmonic nanojunctions, play a crucial role in achieving extreme confinement of the electromagnetic field to Ångström scales [38, 39, 37], resulting in a significant enhancement of Raman scattering [40, 41, 42] and ultrahigh spatial resolution [43, 44, 45, 46].
-Yet, the exceptional sensitivity in TERS/SERS has been demonstrated primarily for organic molecules chemisorbed onto the substrate, where the enhancement benefits from chemical effects arising from the molecule–surface (and molecule–tip) interactions [47].
-Therefore, applying TERS/SERS to physisorbed systems has remained challenging.
-Here we demonstrate that LT-TERS can detect hydrogen molecules physisorbed onto an Ag(111) surface with an Ag tip [Fig.
-1(a), inset], allowing the identification of their rotational and vibrational states.
-Furthermore, precise gap-distance control enables us to investigate not only the contribution of the picocavity field to the Raman scattering process but also the influence of tip proximity on the rotational/vibrational energies of the molecules through non-trivial nuclear quantum effects.
-Density functional theory (DFT), path-integral molecular dynamics (PIMD), and quantum vibrational energy calculations suggest that the tip–molecule–surface potential energy profiles are critically affected by the intermolecular and tip–molecular interactions, leading to unusual isotope effects on vibrational frequencies.
-The LT-TERS experiments were performed in an ultra-high vacuum chamber [see Sect.
-I-A of the Supplemental Material (SM) [48]].
-The 532-nm incident laser beam was linearly polarized along the tip axis and focused on the apex of a focused-ion-beam-sharpened Ag tip [40, 42] using an in-situ parabolic mirror, and the scattering light was detected by an ex-situ spectrometer.
-We conducted DFT calculations [49, 50] of H2/D2 on a Ag(111) (periodic) surface, employing the Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional and screened pairwise vdW interactions [51] (see SM Sect.
+Although the detection demands exceptional sensitivity due to the intrinsically small cross section, single-molecule detection can be achieved by tip-enhanced Raman spectroscopy (TERS) and surface-enhanced Raman spectroscopy (SERS) [33, 34, 35, 36, 37].
+Recent studies have established that picocavities, atomic-scale protrusions present in plasmonic nanojunctions, play a crucial role in achieving extreme confinement of the electromagnetic field to angstrom scales [38, 39, 37], resulting in a significant enhancement of Raman scattering [40, 41, 42] and ultrahigh spatial resolution [43, 44, 45, 46].
+Yet, the exceptional sensitivity in TERS and SERS has been demonstrated primarily for organic molecules chemisorbed onto the substrate, where the enhancement benefits from chemical effects arising from the molecule–surface (and molecule–tip) interactions [47].
+Therefore, applying TERS and SERS to physisorbed systems has remained challenging.
+Here, we demonstrate that low-temperature TERS (LT-TERS) can detect hydrogen molecules physisorbed onto an Ag(111) surface with an Ag tip [Fig.
+1(a), inset], allowing for the identification of their rotational and vibrational states.
+Furthermore, precise gap-distance control enables us to investigate not only the contribution of the picocavity field to the Raman scattering process but also the influence of tip proximity on the rotational and vibrational energies of the molecules through nontrivial nuclear quantum effects.
+Density functional theory (DFT), path-integral molecular dynamics, and quantum vibrational energy calculations suggest that the tip–molecule–surface potential energy profiles are critically affected by the intermolecular and tip–molecular interactions, leading to unusual isotope effects on vibrational frequencies.
+The LT-TERS experiments were performed in an ultrahigh vacuum chamber [see Sec.
+I-A of Supplemental Material (SM) [48]].
+The 532-nm incident laser beam was linearly polarized along the tip axis and focused on the apex of a focused-ion-beam-sharpened Ag tip [40, 42] using an in situ parabolic mirror, and the scattering light was detected by an ex situ spectrometer.
+We conducted DFT calculations [50, 51] of H2/D2 on a Ag(111) (periodic) surface, employing the Perdew-Burke-Ernzerhof exchange-correlation functional and screened pairwise van der Waals interactions [52] (see SM Sec.
 I-B [48]).
-The preferred adsorption sites (fcc/hcp hollow) and binding energies (30 meV) were in very good agreement with previous studies of isolated molecules on the surface [52, 53].
-Unless specified otherwise, we considered a coverage of 0.69 monolayer (ML), motivated by the results of PIMD simulations at LT (see SM Sect.
+The preferred adsorption sites (fcc or hcp hollow) and binding energies (30 meV) were in very good agreement with previous studies of isolated molecules on the surface [55, 65].
+Unless specified otherwise, we considered a coverage of 0.69 monolayer (ML), motivated by the results of path-integral molecular dynamics simulations at low temperature (see SM Sec.
 I-B [48]).
 At 0.69 ML coverage, the intermolecular attractive interaction increases the binding energy by 55 meV compared to the lowest coverage of 0.06 ML considered here (Fig.
-S4).
-Over the sample, we placed the Ag tip model shown in SM Sect.
+S3).
+Over the sample, we placed the Ag-tip model shown in SM Sec.
 I-B [48] (see also Fig.
-S5), which was also previously discussed in Refs.
+S4), which was also previously discussed in Refs.
 [46, 54, 41].
-H2 or D2 gas was introduced into the chamber where a clean Ag(111) surface was cooled at 10 K (see SM Sect.
+H2 or D2 gas was introduced into the chamber where a clean Ag(111) surface was cooled at 10 K (see SM Sec.
 I-A [48]).
-Regardless of laser illumination, no static hydrogen molecule on the surface was imaged with LT-STM (Fig.
+Regardless of laser illumination, no static hydrogen molecules on the surface were imaged with STM (Fig.
 S1).
@@
 Similarly, for D2, intense peaks appeared at 169 and 2967 cm-1, with a weak peak at 3133 cm-1.
-The three peaks can be assigned to rotational/vibrotational modes of hydrogen.
-Note that broad or asymmetric peak shapes are artifacts in the wide wavenumber range spectra (see SM Sect.
+The three peaks can be assigned to rotational and vibrotational modes of hydrogen.
+Note that broad or asymmetric peak shapes are artifacts in the wide wavenumber range spectra (see SM Sec.
 I-A [48]).
@@
 S2).
-Figure 1(b) depicts the energy diagram of rotational/vibrational levels and allowed transitions under the Raman selection rule.
+Figure 1(b) depicts the energy diagram of rotational and vibrational levels and allowed transitions under the Raman selection rule.
 The observed peaks, in order from lowest to highest, correspond to the lowest-energy rotational transition [; “rot” labeled in Fig.
 1(a)], the lowest-energy vibrational transition [; “vib”], and their combination [; “rot+vib”], respectively.
-The three modes were also observed in high-resolution electron energy loss spectroscopy (HREELS) of hydrogen on a polycrystalline Ag film [8] and a Cu(001) surface [9] at approximately 10–12 K; however, better energy resolution (10 cm-1) in TERS than that of the HREELS (28 cm-1 [9]) makes the red-shifts from gas-phase hydrogen more discernible (Table S4).
-We should also note that only the stable nuclear spin isomer of each isotope, i.e., para-H2 and ortho-D2 [Fig.
+The three modes were also observed in high-resolution electron energy loss spectroscopy of hydrogen on a polycrystalline Ag film [8] and a Cu(001) surface [9] at approximately 10–12 K; however, better energy resolution (10 cm-1) in TERS than that of the electron energy loss spectroscopy (28 cm-1 [9]) makes the redshifts from gas-phase hydrogen more discernible (Table S4).
+We should also note that only the stable nuclear-spin isomer of each isotope, i.e., para-H2 and ortho-D2 [Fig.
 1(b)], is detected in TERS (Fig.
 S2).
-On the surface, the stable isomer is dominant due to the ortho–para conversion [8, 4, 52] (see SM Sect.
+On the surface, the stable isomer is dominant due to the ortho–para conversion [8, 4, 55] (see SM Sec.
 II-A [48]).
 The rotational transition at 351 cm-1 for H2/Ag(111) [Fig.
-1(a)] is slightly red-shifted from 354 cm-1 in the gas phase [56].
-We calculate the H–H bonding distance to be 75.4 pm on Ag(111) (cf.
+1(a)] is slightly redshifted from 354 cm-1 in the gas phase [62].
+We calculate the H-H bond length to be 75.4 pm on Ag(111) (cf.
 75.1 pm in the gas phase).
-Considering a rigid-rotor approximation, the rotational transition energy of para-H2 is calculated to be 353 cm-1 on the surface, compared to 356 cm-1 in the gas phase.
-This is in good agreement with the experiment.
-Notably, the H–H bonding distance of H2 on Ag(111) varied by only 0.06 pm at the closest tip position with respect to simulations without any tip.
+Considering a rigid-rotor approximation, the rotational transition energy of para-H2 is calculated to be 353 cm-1 on the surface, compared to 356 cm-1 in the gas phase, showing good agreement with the experiment.
+Notably, the H-H bonding distance of H2 on Ag(111) varied by only 0.06 pm at the closest tip position with respect to simulations without any tip.
 This suggests that the hydrogen molecules on the surface and in the STM junction behave as nearly free rotors.
 In the TERS spectrum [Fig.
-1(a)], the vibrational frequency of 4121 cm-1 exhibits a red-shifted relative to the gas-phase value of 4161 cm-1 [56].
-The vibrational transition energy is calculated to be 4088 cm-1 on the surface (without the Ag tip), after scaling the calculations performed in the harmonic approximation (see SM Sect.
+1(a)], the vibrational frequency of 4121 cm-1 exhibits a redshift relative to the gas-phase value of 4161 cm-1 [62].
+The vibrational transition energy is calculated to be 4088 cm-1 on the surface (without the Ag tip), after applying a scale factor to the calculations performed in the harmonic approximation (see SM Sec.
 I-C [48]; see also Fig.
-S9 for a simulated LT-TERS spectrum [46]).
-The slight discrepancy between the experimental and calculated frequencies are consistent with the typical spurious softening of the PBE functional.
-Compared to the calculated gas-phase value, the vibrational frequency shows a red-shift of 73 cm-1 upon the adsorption, which is larger than, but in reasonable agreement with the experiment.
+S8 for a simulated LT-TERS spectrum [46]).
+The slight discrepancy between the experimental and calculated frequencies is consistent with the typical spurious softening of the Perdew-Burke-Ernzerhof functional.
+Compared to the calculated gas-phase value, the vibrational frequency shows a redshift of 73 cm-1 upon the adsorption, which is larger than, but in reasonable agreement with the experiment.
 The TERS intensity is strongly correlated with the localized surface plasmonic resonance (LSPR) of the tip–sample junction.
-In general, two types of enhancement mechanisms contribute to TERS [36, 47]: electromagnetic enhancement through LSPR and chemical enhancement originating from charge transfer between the target molecule and the tip/surface.
+In general, two types of enhancement mechanisms contribute to TERS [36, 47]: electromagnetic enhancement through LSPR and chemical enhancement originating from charge transfer between the target molecule and the tip or surface.
 In physisorption systems, the electromagnetic enhancement is expected to be dominant.
 To elucidate the role of electromagnetic enhancement, it is crucial to investigate the gap-distance dependence of TERS signals.
-Figure 2(a) shows the waterfall plot of the gap-distance dependent TERS spectra recorded for D2/Ag(111).
+Figure 2(a) shows the waterfall plot of the gap-distance-dependent TERS spectra recorded for D2/Ag(111).
 When the tip approaches the surface, the TERS intensities of the rotational and vibrational modes increase apparently exponentially [Fig.
-2(b)], in agreement with the prediction based on the electric field of a light-induced dipole in the plasmonic tip [57].
-A broad background with a maximum at 590 nm is attributed to electronic Raman scattering which represents the spectral response of the LSPR in the junction [58].
-The gap-distance dependent intensities of both rotational and vibrational modes closely follow the plasmonic feature [Fig.
+2(b)], in agreement with the previously reported analytical model simplifying a plasmonic tip as a point dipole [60].
+A broad background with a maximum at 590 nm is attributed to electronic Raman scattering which reflects the spectral response of the LSPR in the junction [67].
+The gap-distance-dependent intensities of both rotational and vibrational modes closely follow the plasmonic feature [Fig.
 2(b)], indicating the dominant contribution of the gap-mode plasmon to the TERS enhancement.
 Additionally, we observed that changes in both the plasmon background and the molecular Raman peak occurred by an accidental modification of the tip-apex structure (Fig.
-S10), implying that the TERS enhancement is susceptible to the atomic-scale structure of the tip apex, which generates a plasmonic picocavity.
-The gap-distance dependence of the TERS intensities is well reproduced by simulating the electronic field distribution in the Ag-tip–Ag-surface junction using the finite element method (FEM) (see SM [48] for details).
-To model the plasmonic picocavity, we assume an atomic-scale protrusion (0.5-nm radius of curvature) attached to a blunt tip body [40].
+S11), implying that the TERS enhancement is susceptible to the atomic-scale structure of the tip apex, which generates a plasmonic picocavity.
+The gap-distance dependence of the TERS intensities is well reproduced by simulating the electronic field distribution in the Ag-tip–Ag-surface junction using the finite element method (see SM Sec.
+I-E [48]).
+To model the plasmonic picocavity, we assume an atomic-scale protrusion (0.5-nm radius of curvature) attached to a blunt tip body (Fig.
+S9).
 The black curve in Fig.
 2(c) shows the field enhancement at the expected position of a D2 molecule located beneath the Ag tip, indicating a steep increase as the tip approaches.
-We confirmed the critical role of this sub-nanometric structure in TERS; as shown by the orange curve in Fig.
-2(c), the tip without the atomic-scale protrusion exhibits a more moderate increase of the field enhancement as the tip approaches, which does not fit the experimental data [Fig.
+We confirmed the critical role of this subnanometric structure in TERS; as shown by the orange curve in Fig.
+2(c), the tip without the atomic-scale protrusion exhibits a more moderate increase of the field enhancement as the tip approaches, which does not reproduce the experimental data [Fig.
 2(b)].
@@
 2(c), inset], indicating much lower enhancement.
-Therefore, TERS can achieve ‘nearly’ single-molecule sensitivity, even for physisorption systems.
-The contribution of LSPR to the TERS intensity was further confirmed by our excitation-wavelength dependent measurements (SM Sect.
+Therefore, TERS can achieve “nearly” single-molecule sensitivity, even for physisorption systems.
+The contribution of LSPR to the TERS intensity was further confirmed by our excitation-wavelength dependent measurements (SM Sec.
 II-B [48]); a strong TERS peak was observed only when the scattering light wavelength matches with the resonance range of the LSPR (Fig.
-S3).
-TERS reveals the sensitivity of the rotational/vibrational states to the interactions between a target molecule and its local environment.
-Figures 3(a)–3(d) show the gap-distance dependence of high-resolution TERS spectra focusing on the rotational/vibrational modes of H2/D2.
-Notably, only the vibrational frequency of H2 exhibits a significant red-shift as the gap distance decreases [Fig.
+S10).
+TERS reveals the sensitivity of the rotational and vibrational states to the interactions between a target molecule and its local environment.
+Figures 3(a)–3(d) show the gap-distance dependence of high-resolution TERS spectra focusing on the rotational and vibrational modes of H2 and D2.
+Notably, only the vibrational frequency of H2 exhibits a significant redshift as the gap distance decreases [Fig.
 3(c)]; H2 shows a peak shift of 33 cm-1 with a 200 pm displacement of the tip, whereas is merely cm-1 for D2.
-Additionally, the peak width of H2 vibrational mode also increases as the gap distance decreases (see SM Sect.
+Additionally, the peak width of the H2 vibrational mode also increases as the gap distance decreases (see SM Sec.
 II-C [48] for details).
@@
 3(a)] and D2 [3(b)] are almost unchanged, consistent with our calculations showing that the interatomic distance is independent of the tip height, as described above.
-The isotope dependence of the red-shift of the vibrational band far exceeds the scale of typical isotope effects.
-To understand this notable behavior, we calculated the potential energy profile along the rigid vertical displacement of the center of mass of one molecule between the surface and the tip apex [Fig.
+The isotope dependence of the redshift of the vibrational band far exceeds the scale of typical isotope effects.
+To understand this anomalous behavior, we calculated the potential energy profile along the rigid vertical displacement of the center of mass of one molecule between the surface and the tip apex [Fig.
 4(a)], while varying the tip–sample distance [see profiles (i)–(iv) at different in Fig.
 4(b)].
-At each position, we calculated the H2 and D2 harmonic intramolecular vibrational frequency (see SM Sect.
-II-C [48]).
+At each position, we calculated the H2 and D2 harmonic intramolecular vibrational frequency (see SM Sec.
+I-C [48]).
 The calculations indicate a pronounced coupling of the stretch coordinate with the molecular translation along , evidenced by a steep variation of along this coordinate (Fig.
-S8).
+S7).
 Furthermore, this variation changes depending on the tip–sample distance, leading us to define .
-As detailed in SM Sect.
+As detailed in SM Sec.
 I-C [48], for each , we constructed a model 2D Schrödinger equation in the coupled coordinates and , parameterized by and .
 The solution of this model gives us the expected intramolecular stretch vibrational transition energy for each isotope, fully accounting for the anharmonic coupling between these coordinates.
-We concluded that if both H2 and D2 probe the same well of [Fig.
-4(b)], the isotope effect on can account at most for a factor 2 between H2 and D2 (Tables S2 and S3), in disagreement with the experiment (Fig.
+We concluded that if both H2 and D2 probe the same potential energy profile [Fig.
+4(b)], for reasonable position-dependent stretch frequencies the calculated peak shifts are in a ratio of at most 2:1 for H2:D2 (Tables S2 and S3), in disagreement with the measured 16:1 ratio (Fig.
 3).
 We demonstrate that this discrepancy can be resolved by accounting for density differences between H2 and D2.
-At lower temperatures, the solids formed by H2 and D2 have different molar volumes due to nuclear quantum effects [zero-point energy (ZPE) swelling [59]; see also SM Sect.
+At lower temperatures, the solids formed by H2 and D2 have different molar volumes due to nuclear quantum effects [zero-point energy (ZPE) swelling [68]; see also SM Sec.
 I-B [48]].
@@
 Consequently, the average position of the molecule center of mass [bullet points in Figs.
-4(b) and (d)] also shifts towards the tip as decreases [dotted line in Fig.
+4(b) and (d)] also shifts toward the tip as decreases [dotted line in Fig.
 4(d)].
@@
 As shown in Fig.
-4(e), considering the higher-density scenario for D2 (the filled bullets) and the lower-density scenario for H2 (the empty bullets) can account for the large isotope effect in the red-shift of the stretch transition as the tip approaches.
+4(e), considering the higher-density scenario for D2 (the filled bullets) and the lower-density scenario for H2 (the empty bullets) can account for the large isotope effect in the redshift of the stretch transition as the tip approaches.
 Figure 4(e) also shows the experimental values from the plots in Fig.
 3 (the solid and dotted curves for H2 and D2, respectively; see also Fig.
-S11).
-Although the exact gap distance in STM experiment is hard to determine (see SM Sect.
+S12).
+Although the exact gap distance in STM experiment is hard to determine (see SM Sec.
 II-C [48]), the overlap of the experimental curves with the calculated plots in Fig.
 4(e) shows good agreement.
-The remaining numerical discrepancies can be attributed to the limitation of the DFT level of theory (see SM Sect.
-I-B [48]), the exact tip apex geometry, and to the uncertainty in determining the exact local surface coverage.
+The remaining numerical discrepancies can be attributed to the limitation of the DFT level of theory (see SM Sec.
+I-B [48]), the exact tip-apex geometry, and to the uncertainty in determining the exact local surface coverage.
 Nevertheless, our results highlight the crucial role of intermolecular interactions in accurately interpreting vibrational spectroscopy of hydrogen isotopes in nanojunctions, where nuclear quantum effects have a significant impact [13].
-In summary, with LT-TERS, we observed the rotational/vibrational transitions of local H2 and D2 molecules through the interaction with the picocavity field.
-Our results show that LT-TERS is applicable to weakly adsorbed molecules, offering deeper insights into the local structures, reactions, and dynamics of adsorbates, such as chemical reactions at active sites [61] and surface diffusion (spillover) in catalytic processes [62], at the single-molecule level.
+In summary, with LT-TERS, we observed the rotational and vibrational transitions of local H2 and D2 molecules through the interaction with the picocavity field.
+Our results demonstrate that LT-TERS is applicable to weakly adsorbed molecules, offering deeper insights into the local structures, reactions, and dynamics of adsorbates, such as chemical reactivity at active sites [70] and surface diffusion (spillover) in catalytic processes [71], at the single-molecule level.
 Acknowledgements.
-We thank Adnan Hammud for the Ag-tip fabrication.
+Acknowledgments—We thank Adnan Hammud for the Ag-tip fabrication.
 We thank Jun Yoshinobu and Heiko Appel for helpful discussions.
@@
 acknowledges the support of JST FOREST Grant No.
-JPMJFR201J and JSPS KAKENHI Grant No.
-JP19K24684.References
+JPMJFR201J.References
 - Ockwig and Nenoff [2007] N. W. Ockwig and T. M. Nenoff, Membranes for hydrogen separation, Chem. Rev. 107, 4078 (2007).
@@
 - Vidali et al. [1991] G. Vidali, G. Ihm, H.-Y. Kim, and M. W. Cole, Potentials of physical adsorption, Surf. Sci. Rep. 12, 135 (1991).
-- Bruch et al. [2007] L. Bruch, R. Diehl, and J. Venables, Progress in the measurement and modeling of physisorbed layers, Rev. Mod. Phys. 79, 1381 (2007).
-- Avouris et al. [1982] P. Avouris, D. Schmeisser, and J. Demuth, Observation of rotational excitations of H2 adsorbed on Ag surfaces, Phys. Rev. Lett. 48, 199 (1982).
+- Bruch et al. [2007] L. W. Bruch, R. D. Diehl, and J. A. Venables, Progress in the measurement and modeling of physisorbed layers, Rev. Mod. Phys. 79, 1381 (2007).
+- Avouris et al. [1982] P. Avouris, D. Schmeisser, and J. E. Demuth, Observation of rotational excitations of H2 adsorbed on Ag surfaces, Phys. Rev. Lett. 48, 199 (1982).
 - Andersson and Harris [1982] S. Andersson and J. Harris, Observation of rotational transitions for H2, D2, and HD adsorbed on Cu(100), Phys. Rev. Lett. 48, 545 (1982).
-- Yu et al. [1983] C.-f. Yu, K. B. Whaley, C. S. Hogg, and S. J. Sibener, Selective adsorption resonances in the scattering of -H2, -H2, -D2 and -D2 from Ag(111), Phys. Rev. Lett. 51, 2210 (1983).
+- Yu et al. [1983] C. F. Yu, K. B. Whaley, C. S. Hogg, and S. J. Sibener, Selective adsorption resonances in the scattering of -H2, -H2, -D2 and -D2 from Ag(111), Phys. Rev. Lett. 51, 2210 (1983).
 - Sakurai et al. [1988] M. Sakurai, T. Okano, and Y. Tuzi, Ortho-para conversion of -H2 physisorbed on Ag(111) near two-dimensional condensation conditions, Appl. Surf. Sci. 33, 245 (1988).
@@
 Wilde, W.
+A.
 Diño, M.
@@
 - Sugimoto et al. [2017] T. Sugimoto, Y. Kunisada, and K. Fukutani, Inelastic electron tunneling mediated by a molecular quantum rotator, Phys. Rev. B 96, 241409(R) (2017).
-- Gupta et al. [2005] J. Gupta, C. Lutz, A. Heinrich, and D. Eigler, Strongly coverage-dependent excitations of adsorbed molecular hydrogen, Phys. Rev. B 71, 115416 (2005).
+- Gupta et al. [2005] J. A. Gupta, C. P. Lutz, A. J. Heinrich, and D. M. Eigler, Strongly coverage-dependent excitations of adsorbed molecular hydrogen, Phys. Rev. B 71, 115416 (2005).
 - Temirov et al.
@@
 - Wang et al. [2023] L. Wang, D. Bai, Y. Xia, and W. Ho, Electrical manipulation of quantum coherence in a two-level molecular system, Phys. Rev. Lett. 130, 096201 (2023).
-- Thijssen et al. [2006] W. Thijssen, D. Djukic, A. Otte, R. Bremmer, and J. van Ruitenbeek, Vibrationally induced two-level systems in single-molecule junctions, Phys. Rev. Lett. 97, 226806 (2006).
+- Thijssen et al.
+[2006] W.
+H.
+A.
+Thijssen, D.
+Djukic, A.
+F.
+Otte, R.
+H.
+Bremmer, and J.
+M.
+van Ruitenbeek, Vibrationally induced two-level systems in single-molecule junctions, Phys.
+Rev.
+Lett.
+97, 226806 (2006).
 - Halbritter et al. [2008] A. Halbritter, P. Makk, S. Csonka, and G. Mihály, Huge negative differential conductance in Au–H2 molecular nanojunctions, Phys. Rev. B 77, 075402 (2008).
 - Li et al. [2015] Y. Li, S. Kaneko, S. Fujii, and M. Kiguchi, Symmetry of single hydrogen molecular junction with Au, Ag, and Cu electrodes, J. Phys. Chem. C 119, 19143 (2015).
-- Trouwborst et al. [2009] M. Trouwborst, E. Huisman, S. van der Molen, and B. van Wees, Bistable hysteresis and resistance switching in hydrogen-gold junctions, Phys. Rev. B 80, 081407(R) (2009).
+- Trouwborst et al.
+[2009] M.
+L.
+Trouwborst, E.
+H.
+Huisman, S.
+J.
+van der Molen, and B.
+J.
+van Wees, Bistable hysteresis and resistance switching in hydrogen-gold junctions, Phys.
+Rev.
+B 80, 081407(R) (2009).
 - Haran [2010] G. Haran, Single-molecule Raman spectroscopy: a probe of surface dynamics and plasmonic fields, Acc. Chem. Res. 43, 1135 (2010).
@@
 [49, 40, 50, 51, 52, 53, 46, 54, 41, 55, 56, 17, 57, 58, 59, 60, 61, 4, 62, 8, 11, 14, 15, 9, 63, 64, 18].
+- Summers [1969] R. L. Summers, Empirical observations on the sensitivity of hot cathode ionization type vacuum gages, NASA Technical Report NASA-TN-D-5285 (1969).
 - Blum et al.
@@
 108, 146103 (2012).
-- Kunisada and Kasai [2015] Y. Kunisada and H. Kasai, Hindered rotational physisorption states of H2 on Ag(111) surfaces, Phys. Chem. Chem. Phys. 17, 19625 (2015).
-- Smeets and Kroes [2021] E.
-W.
-Smeets and G.-J.
-Kroes, Designing new SRP density functionals including non-local vdW-DF2 correlation for H2 + Cu(111) and their transferability to H2 + Ag(111), Au(111) and Pt(111), Phys.
-Chem.
-Chem.
-Phys.
-23, 7875 (2021).
+- Hermann and Tkatchenko [2020] J.
+Hermann and A.
+Tkatchenko, Density functional model for van der Waals interactions: Unifying many-body atomic approaches with nonlocal functionals, Phys.
+Rev.
+Lett.
+124, 146401 (2020).
 - Cirera et al.
@@
 22, 2170 (2022).
-- Cui et al.
-[2023] J.
-Cui, Y.
-Zhang, and Z.-C.
-Dong, Atomistic polarization model for Raman scattering simulations of large metal tips with atomic-scale protrusions at the tip apex, Nanophotonics 12, 4031 (2023).
-- Veirs and Rosenblatt [1987] D. K. Veirs and G. M. Rosenblatt, Raman line positions in molecular hydrogen: H2, HD, HT, D2, DT, and T2, J. Mol. Spectrosc. 121, 401 (1987).
-- Pettinger et al.
-[2009] B.
-Pettinger, K.
-F.
-Domke, D.
-Zhang, G.
-Picardi, and R.
-Schuster, Tip-enhanced Raman scattering: influence of the tip-surface geometry on optical resonance and enhancement, Surf.
-Sci.
-603, 1335 (2009).
-- Kamimura et al. [2022] R. Kamimura, T. Kondo, K. Motobayashi, and K. Ikeda, Surface-enhanced electronic Raman scattering at various metal surfaces, Phys. Status Solidi B 259, 2100589 (2022).
-- Markland et al. [2008] T. E. Markland, S. Habershon, and D. E. Manolopoulos, Quantum diffusion of hydrogen and muonium atoms in liquid water and hexagonal ice, J. Chem. Phys. 128, 194506 (2008).
-- Megaw and Simon [1936] H. D. Megaw and F. Simon, Density and compressibility of solid hydrogen and deuterium at 4.2 K, Nature 138, 244 (1936).
-- Cai et al. [2023] Z.-F. Cai, N. Kumar, and R. Zenobi, Probing on-surface chemistry at the nanoscale using tip-enhanced Raman spectroscopy, CCS Chem. 5, 55 (2023).
-- Yang et al. [2024] J.-L. Yang, Y.-W. Zhou, M.-Y. Jia, Q. Liu, and Z.-F. Cai, Insights into heterogeneous catalysis by tip-enhanced Raman spectroscopy, ChemCatChem 16, e202400370 (2024).
-- Summers [1969] R. L. Summers, Empirical observations on the sensitivity of hot cathode ionization type vacuum gages, NASA Technical Report NASA-TN-D-5285 (1969).
-- Hermann and Tkatchenko [2020] J.
-Hermann and A.
-Tkatchenko, Density functional model for van der Waals interactions: Unifying many-body atomic approaches with nonlocal functionals, Phys.
-Rev.
-Lett.
-124, 146401 (2020).
+- Kunisada and Kasai [2015] Y. Kunisada and H. Kasai, Hindered rotational physisorption states of H2 on Ag(111) surfaces, Phys. Chem. Chem. Phys. 17, 19625 (2015).
 - Kapil et al.
@@
 - Jiang et al. [2016] Y. Jiang, S. Pillai, and M. A. Green, Realistic silver optical constants for plasmonics, Sci. Rep. 6, 30605 (2016).
+- Pettinger et al.
+[2009] B.
+Pettinger, K.
+F.
+Domke, D.
+Zhang, G.
+Picardi, and R.
+Schuster, Tip-enhanced Raman scattering: influence of the tip-surface geometry on optical resonance and enhancement, Surf.
+Sci.
+603, 1335 (2009).
 - Wolniewicz [1983] L. Wolniewicz, The state vibration-rotational energies of the H2, HD, and D2 molecules, J. Chem. Phys. 78, 6173 (1983).
+- Veirs and Rosenblatt [1987] D. K. Veirs and G. M. Rosenblatt, Raman line positions in molecular hydrogen: H2, HD, HT, D2, DT, and T2, J. Mol. Spectrosc. 121, 401 (1987).
 - Teal and MacWood [1935] G. K. Teal and G. E. MacWood, The Raman spectra of the isotopic molecules H2, HD, and D2, J. Chem. Phys. 3, 760 (1935).
-- Persson and Ryberg [1989] B. Persson and R. Ryberg, Vibrational line shapes of low-frequency adsorbate modes: CO on Pt(111), Phys. Rev. B 40, 10273 (1989).
+- Persson and Ryberg [1989] B. N. J. Persson and R. Ryberg, Vibrational line shapes of low-frequency adsorbate modes: CO on Pt(111), Phys. Rev. B 40, 10273 (1989).
+- Smeets and Kroes [2021] E.
+W.
+Smeets and G.-J.
+Kroes, Designing new SRP density functionals including non-local vdW-DF2 correlation for H2 + Cu(111) and their transferability to H2 + Ag(111), Au(111) and Pt(111), Phys.
+Chem.
+Chem.
+Phys.
+23, 7875 (2021).
+- Cui et al.
+[2023] J.
+Cui, Y.
+Zhang, and Z.-C.
+Dong, Atomistic polarization model for Raman scattering simulations of large metal tips with atomic-scale protrusions at the tip apex, Nanophotonics 12, 4031 (2023).
+- Kamimura et al. [2022] R. Kamimura, T. Kondo, K. Motobayashi, and K. Ikeda, Surface-enhanced electronic Raman scattering at various metal surfaces, Phys. Status Solidi B 259, 2100589 (2022).
+- Markland et al. [2008] T. E. Markland, S. Habershon, and D. E. Manolopoulos, Quantum diffusion of hydrogen and muonium atoms in liquid water and hexagonal ice, J. Chem. Phys. 128, 194506 (2008).
+- Megaw and Simon [1936] H. D. Megaw and F. Simon, Density and compressibility of solid hydrogen and deuterium at 4.2 K, Nature 138, 244 (1936).
+- Cai et al. [2023] Z.-F. Cai, N. Kumar, and R. Zenobi, Probing on-surface chemistry at the nanoscale using tip-enhanced Raman spectroscopy, CCS Chem. 5, 55 (2023).
+- Yang et al. [2024] J.-L. Yang, Y.-W. Zhou, M.-Y. Jia, Q. Liu, and Z.-F. Cai, Insights into heterogeneous catalysis by tip-enhanced Raman spectroscopy, ChemCatChem 16, e202400370 (2024).
 Supplemental Material: Picocavity-Enhanced Raman Spectroscopy of Physisorbed H2 and D2 Molecules
@@
 I.1 A. STM-TERS Experiments
-The experiments were performed in an ultrahigh vacuum (UHV) chamber equipped with an STM-based TERS machine (modified UNISOKU USM-1400) at a sample temperature of 10 K.
+The scanning tunneling microscopy (STM) and tip-enhanced Raman spectroscopy (TERS) experiments were performed in an ultrahigh vacuum (UHV) chamber (modified UNISOKU USM-1400) at a sample temperature of 10 K.
 A single-crystalline Ag(111) surface (MaTeck) was cleaned by Ar-sputtering-and-annealing cycles in the chamber.
@@
 The bias voltage was applied to the sample while the tip was grounded.
-The STM images were obtained in the constant current mode at = 50 pA.
+The STM images were obtained in the constant current mode at tunneling current of 50 pA.
 The cleanness of the Ag(111) surface and tip was confirmed by STM images [Fig.
@@
 S2, at 0 L).
-In the STM junction, -polarised visible laser (532-nm solid-state or 633-nm HeNe) was directed through a fused silica window and an Ag-coated parabolic mirror (numerical aperture of 0.6).
-The incident laser power is 5.9 mW at = 532 nm and 9.7 mW at 633 nm.
-The parabolic mirror is mounted on a five-axis piezoelectric-motor stage in the LT-STM unit and the laser is focused at the tip apex by adjusting the position and angle of the parabolic mirror.
+In the STM junction, -polarized visible laser (532-nm solid-state or 633-nm HeNe) was directed through a fused silica window and an Ag-coated parabolic mirror.
+The parabolic mirror is mounted on a five-axis piezoelectric-motor stage in the low-temperature (LT) STM unit and the laser is focused at the tip apex by adjusting the position and angle of the parabolic mirror.
 The scattered light collected by the parabolic mirror was directed to a spectrometer (AndorShamrock 303i) outside the UHV chamber via a beam splitter and a longpass filter, each selected according to the excitation wavelength .
+The nominal incident laser power is 5.9 mW at = 532 nm and 9.7 mW at = 633 nm.
 For the TERS and STM-induced luminescence (STML) spectra, we used the gratings 150 (for Fig.
-S3), 600 (Figs.
+S10), 600 (Figs.
 1, 2, and S2), and 1200 line/mm (Fig.
-3 and S10).
+3 and S11).
 A lower-density grating allows the acquisition of spectra over a wider wavelength range, as shown in Fig.
-S3 (see Sect.
+S10 (see Sec.
 II-B), but resulting in broad and asymmetric Raman peaks.
@@
 I.2 B. DFT Calculations and Molecular Dynamics
-DFT calculations were performed with the FHI-aims program package [49, 50].
+Density functional theory (DFT) calculations were performed with the FHI-aims program package [50, 51].
 We used intermediate defaults for basis sets, radial grids and cutoff potential parameters and light defaults for the Hartree multipole expansion and angular grids.
@@
 We employed a -point grid and a dipole correction was employed to maximally decouple periodic images along the direction.
-We used the PBE exchange-correlation functional and included vdW interactions through the pairwise screened TS-vdWsurf method of Ref.
+We used the Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional and included van der Waals (vdW) interactions through the pairwise screened TS-vdWsurf method of Ref.
 [52] in all simulations reported in the main text.
-Figure S4 shows a comparison of binding energies with the many-body MBD-NL dispersion correction of Ref.
+Figure S3 shows a comparison of binding energies with the many-body dispersion correction termed MBD-NL in Ref.
 [53], indicating a weaker binding energy by only 5 meV and a binding distance 0.1 Å closer to the surface.
@@
 The geometry of the tip is shown in Fig.
-S5.
+S4.
 Although this tip size would not be sufficient to capture a proper plasmonic response of the tip, the tip structure can capture the local interactions with molecules close to the tip apex [46, 54, 41].
@@
 The blue curve in Fig.
-S4 shows the binding curve of a single H2 molecule adsorbed on the fcc-hollow site of the Ag(111) surface, indicating an equilibrium H2–Ag(111) bonding length of 3.09 Å and a binding energy of 32 meV.
+S3 shows the binding curve of a single H2 molecule adsorbed on the face-centered-cubic (fcc) hollow site of the Ag(111) surface, indicating an equilibrium H2–Ag(111) bonding length of 3.09 Å and a binding energy of 32 meV.
 The model we used in most simulations contained 11 molecules in the unit cell (see rationale below).
 For this molecular concentration, we evaluated the intermolecular interaction by comparing the binding curves (orange curve in Fig.
-S4) with that of the single-molecule case (blue); the molecule–molecule interaction in our models amounts to 55 meV/molecule.
+S3) with that of the single-molecule case (blue); the molecule–molecule interaction in our models amounts to 55 meV/molecule.
 Performing vibrational analysis on the optimized single molecule adsorbed on the surface showed that negative frequencies related to rotational and translational modes of the molecule were always present.
 This led us to consider higher coverages to model local stable clustering of the molecules on the surface.
-We considered a full monolayer (1 ML), where all fcc-hollow sites of the unit cell were occupied by H2.
-With this setup we performed Born-Oppenheimer molecular dynamics (BOMD) and PIMD at LT ( = 20 K) with the i-PI code [65] connected to FHI-aims.
+We started by considering a full monolayer (1 ML, i.e., 16 molecules on the surface cell), where all fcc-hollow sites of the unit cell were occupied by H2.
+With this setup we performed Born-Oppenheimer molecular dynamics (BOMD) and path-integral molecular dynamics (PIMD) at LT ( = 20 K) with the i-PI code [56] connected to FHI-aims.
 PIMD simulations used 12 beads in conjunction with a normal-mode colored-noise thermostat with 8 auxiliary variables and .
 This number of beads is not fully converged at this temperature even with the colored-noise quantum thermostats, but the purpose of these simulations was exploratory.
-While the full monolayer was stable during more than 10 ps on the classical trajectories, in the quantum trajectory 5 molecules readily desorbed (within the first picosecond) from the substrate.
+While the full monolayer was stable during more than 10 ps on the classical trajectories, in the quantum trajectory 5 molecules readily desorbed (within the first picosecond) from the substrate, leaving only 11 molecules adsorbed, corresponding to a ML coverage.
+This is the coverage adopted in other simulations in this Letter as the high-density model.
 For the quantum case, we validated this observation by simulating trajectories with a parameterized Lennard-Jones potential representing the Ag(111) surface.
@@
 The profile shapes are similar to an asymmetric double-well potential at closer tip–surface distances, while they assume a shape closer to a Morse potential at larger distances, which asymptotically approaches the profile of an H2 molecule on Ag(111) without the Ag tip (yellow curve in Fig.
-S6).
+S5).
 The asymmetry of the double-well is strongly impacted by the presence of other molecules on the surface (molecule–molecule interactions).
@@
 In that case (orange curve), if the molecule can overcome the barrier, it will have a higher probability of being found closer to the tip, which was predicted in a Au-tip–single-H2–Au-surface junction [17].
-As we discuss in the main text, this effect is essential to explain the isotope-dependent red-shift observed experimentally.
+As we discuss in the main text, this effect is essential to explain the isotope-dependent redshift observed experimentally.
 Figure S6(b) shows a comparison of the 0.69 ML coverage with the tip fixed at Å and different vdW dispersion corrections, added to the PBE functional.
 It is clear that details of the vdW dispersion impact the shape of the potential.
-In this case, the many-body vdW dispersion decreases the distance between the two minima of the asymmetric double-well
+In this case, the many-body vdW dispersion decreases the distance between the two minima of the asymmetric double-well.
+This change would impact the quantitative value of the tip–surface distance at which H2 molecules would prefer to sit closer to the tip, at the low-density limit.
+We do not expect qualitative changes on the effects we discuss in the main text.
 I.3 C. Model 2D Schrödinger Equation
@@
 I.2.
-The vibrational motion is parametrized by the reduced mass and angular frequency .
-We set the mass of a hydrogen atom to and use .
+The vibrational motion along the intramolecular stretching coordinate is parametrized by the reduced mass and angular frequency .
+We set the mass of a hydrogen atom to and use , with denoting the mass of the electron.
 The vibrational frequencies are computed from numerical (finite-difference) second derivatives of the DFT energies with respect to the vibrational coordinate,
@@
 Given the large separation in the timescales of translational and vibrational motion, we also considered the vibrationally adiabatic approximation, whereby one fixes the translational coordinate in Eq.
-(S1) and solves the corresponding one-dimensional (1D) Schrödinger equation in , yielding the analytical harmonic vibrational energies
-These are then used to construct the vibrationally adiabatic translational potential,
+(S1) and solves the corresponding one-dimensional (1D) Schrödinger equation in , with known energy eigenvalues
+where is the reduced Planck constant and is the vibrational quantum number. These are then used to construct the vibrationally adiabatic translational potential,
 for which we solve the 1D Schrödinger equation numerically using DVR.
@@
 Given the accuracy of the vibrationally adiabatic approximation, we adopt it in all subsequent calculations.
+[…]
 2D | ||||
 1D |
-[…]
 11.4 | ||||||
@@
 We do not observe any noteworthy features in the calculated spectrum and it overall agrees with the experimental spectrum.
-The red-shift of the rotational band in the simulation is explained by the absence of nuclear spin in the simulations, making calculations reproduce only classical rotations that are red-shifted compared to the quantized transition of para-H2.
+The redshift of the rotational band in the simulation is explained by the absence of nuclear spin in the simulations, making calculations reproduce only classical rotations.
+These are incapable of capturing the quantization of the rotational transitions and thus do not capture the transition of para-H2, where is the rotational quantum number.
+This explains the redshift observed in the simulated frequency of this transition with respect to the experimental value.
 This simulation is of exploratory character, as achieving statistical convergence through the use of parametrized machine-learned potentials along with accurate plasmonic enhancements is beyond the scope of this paper.
 I.5 E. FEM Simulations
-To evaluate the electric field in the STM junction, we performed FEM simulations using COMSOL Multiphysics (version 6.2 with the Wave Optics Module).
-We used the three-dimentional model with an Ag tip and Ag sample plate described in Supporting Information of Ref.
-[40].
-To reproduce a plasmonic picocavity in the STM junction, at the apex of the Ag tip (a length of 300 nm, an effective radius of 30 nm), an Ag half sphere with a radius of 0.5 nm is attached as an atomic-scale protrusion.
-For comparison, we also constructed another Ag tip model with an effective radius of 30 nm without the atomic-scale protrusion [see the orange dotted curve in Fig.
-2(c) in the main text].
-A -polarized plane wave with of 532 nm and an electric field of 1 V/m is incident with an angle of 55∘ relative to the sample normal.
+To evaluate the electric field in the STM junction, we performed finite element method (FEM) simulations using COMSOL Multiphysics (version 6.2 with the Wave Optics Module).
+We used the three-dimensional model with an Ag tip and Ag sample plate in a vacuum [40].
+The tip shaft is elongated in direction with a half opening angle of 6∘ from a 30-nm radius sphere [Fig.
+S9(a)].
+The tip is axisymmetrical to [see the inset of Fig.
+2(c) for the coordinate] with a length of 300 nm.
+To reproduce a plasmonic picocavity in the STM junction, at the apex of the Ag tip, an Ag half sphere with a radius of 0.5 nm is attached as an atomic-scale protrusion [Fig.
+S9(b)].
+For comparison [the orange dotted curve in Fig.
+2(c) of the main text], we also constructed another Ag tip model with an effective radius of 30 nm without the atomic-scale protrusion.
+A -polarized plane wave with = 532 nm and an electric field of 1 V/m is incident with an angle of 55∘ relative to the axis.
 The perfectly matched layers surrounding the volume are set to absorb all outgoing waves.
 The dielectric constant of the Ag objects is referenced from the literature [59].
-The surface normal component of the electric field is sampled.
-An Ag flat plate is placed beneath the tip apex with a gap distance .
-Here we define the origin of the gap distance ( = 0) as the tip height when the Ag-tip apex forms point contact with the Ag surface.
-As shown in Table S3, the molecular height is calculated to be 0.32–0.33 nm for D2/Ag(111).
-In the definition, taking into account the atomic radius of Ag, the position of the molecule corresponds to 0.2 nm above the Ag surface plane in the FEM model.
-Although the calculated depends on the tip height (Table S3), in the FEM model, we used the constant height for the sampling point at any gap distances.
-We confirmed that differences in the definition or the molecular position on the order of 0.1 nm do not result in qualitative differences in the field enhancement factor curves.
+The component of the electric field is sampled.
+An Ag flat plate with a thickness of 100 nm is placed beneath the tip apex with a gap distance , as shown in [Fig.
+S9(b).
+As shown in Table S3, the molecular height is calculated to be 0.32–0.33 nm for D2/Ag(111) [Fig.
+S9(c)].
+In the definition, taking into account the atomic radius of Ag, the position of the molecule corresponds to 0.2 nm above the Ag surface plane in the FEM model [the open bullet in Fig.
+S9(b)].
+Although the calculated depends on the tip height (Table S3), the FEM model used the constant height for the sampling point at any .
+We confirmed that differences in the molecular position on the order of 0.1 nm do not result in qualitative differences in the curve shape.
 To directly compare with the experimental TERS intensities, the electric field enhancement factor (Ref.
-[57]) is displayed in the plots of Fig.
-2(c) in the main text.
+[60]) is displayed in Fig.
+2(c).
 To the field enhancement of the nearest neighboring molecule, assuming the molecule is placed in the next adsorption site of Ag(111), = 0.29 nm is used [red triangle in the inset of Fig.
-2(c) in the main text].
+2(c)].
 Thus, when the point at the surface just under the tip is defined as (, , ) = (0, 0, 0), the black and red curves shown in Fig.
@@
 Considering the Ag-atom radius, for the simulations (Fig.
-4 in the main text) is comparable to nm.
-Therefore, pm corresponds to nm, which was used for the alignment of the experimental and theoretical plots displayed in Fig.
-4(e) of the main text.
+4 in the main text) is comparable to nm [Fig.
+S9(b) and (c)].
+Therefore, pm corresponds to nm, which was used for the alignment of the experimental and theoretical plots of the vibrational frequency shifts [Fig.
+4(e)].
 II II. Supplemental Results and Discussion
@@
 para-H2 (nuclear spin quantum number = 0) is more stable than ortho-H2 ( = 1) by 14 meV [61].
-para-H2 has only even rotational quantum numbers , whereas and ortho-H2 has only odd .
+para-H2 has only even , whereas and ortho-H2 has only odd .
 Although ortho-H2 is the predominant species in the gas phase at room temperature (a temperature equilibrium composition of 75%), the molecular adsorption on metals enhances the ortho-to-para conversion, leading to the dominance of the para species on the surfaces [4].
@@
 1 in the main text).
-The dominance of para-H2 on Ag(111) in the TERS spectra ( = 10 K) is consistent with a previous HREELS study of H2 on an Ag film (exposures of 0.5–10 L, = 10 K) [8], whereas some previous reports detect ortho-H2 on Ag(111).
+The dominance of para-H2 on Ag(111) in the TERS spectra ( = 10 K) is consistent with a previous study with high-resolution electron energy loss spectroscopy (HREELS) for H2 on an Ag film (exposures of 0.5–10 L, = 10 K) [8], whereas some previous reports detect ortho-H2 on Ag(111).
 A HREELS study on H2/Ag(111) at 6 K [11] shows that rotational mode peaks of both para-H2 and ortho-H2 were detected under a higher H2 pressure ( Torr) while para-H2 was dominantly detected under a lower pressure ( Torr).
@@
 For our LT-TERS measurement, in contrast, only a monolayer would be formed at the higher temperature, 10 K.
-Another possibility that the metastable isomer were not detected in TERS is that the LSPR in the gap is accelerating the para-to-ortho conversion of the molecule under the tip.
+Another possibility that the metastable isomer were not detected in TERS is that the localized surface plasmonic resonance (LSPR) in the gap is accelerating the para-to-ortho conversion of the molecule under the tip.
 Although far-field 532-nm irradiation did not affect the conversion rate [15], the strong electric field localized in the STM junction could interact with the nuclear spin of the adsorbate.
@@
 After recording the STML, we exposed the surface to the D2 gas, conducting TERS measurements with = 532 and 633 nm [Figs.
-S3(b) and S3(c), respectively].
+S10(b) and S10(c), respectively].
 The peak intensity of each mode significantly changes with , whereas the Raman shift value is unchanged.
-The vibrational intensity is weakened with = 633 nm, indicating the corresponding Raman scattering is not efficiently enhanced by the LSPR because the peak position ( = 779 nm) is located in the weak resonance range of the LSPR [Fig.
-S3(a)].
+The vibrational intensity is weakened with = 633 nm, indicating the corresponding Raman scattering is not efficiently enhanced by the LSPR because the peak position (a wavelength of the Raman scattering of 779 nm) is located in the weak resonance range of the LSPR [Fig.
+S10(a)].
 In contrast, both the rotational and vibrational peaks become intense with = 532 nm because they are more resonant with the LSPR.
@@
 Figure S11(b) shows the horizontal line profiles of the waterfall plot for the H2 “vib” mode at the points indicated by the arrows in the same color.
-The three spectra indicates the red-shift and broadening of the H2 “vib” peak as the gap distance decreases.
+The three spectra indicates the redshift and broadening of the H2 “vib” peak as the gap distance decreases.
 During recording the waterfall TERS plot, we occasionally observed a sudden intensity change caused presumably by the modification of the tip apex structure due to the strong tip–surface interaction at the small gap.
@@
 3 in the main text; see also Fig.
-S10(b)] was fitted by a Gaussian peak with a constant background.
+S11(b)] was fitted by a Gaussian peak with a constant background.
 The top and bottom panels of Fig.
-S11 shows the peak position and width, respectively, as a function of the tip height for the rotational (a) and vibrational (b) modes.
-The error bars indicate the standard deviation of the fitting.
+S12 shows the peak position and width, respectively, as a function of the tip height for the rotational (a) and vibrational (b) modes.
+To display the experimental peak-shift values , the reference point should be specified.
+Ideally is the furthest tip height at which the peak was detected.
+However, since using a too weak peak as the reference value causes large errors in the plot, we adopted as the tip height at which the error value below 2 cm-1 based on the peak fitting.
+The error bars in Fig.
+S12 indicate the standard deviation of the fitting.
 As the tip approaches, the peak intensity increases, reducing the error.
-The point at which the error becomes below 2 cm-1 by the tip approach was used as the reference point [the arrows in the top panel of Fig.
-S11(a) and (b)] and the peak position at this reference point was set to the origin of the peak shift (0 cm-1 of the top panel).
+As shown in the arrows in Fig.
+S12, we set , , , pm for H2 rot, D2 rot, H2 vib, and D2 vib, respectively, determining the origin of each peak shift.
 For the H2 and D2 “vib” modes [top panel of Fig.
-S11(b)], the plots after the reference points were used as the experimental plots shown in Fig.
+S12(b)], the plots at were used as the experimental curves shown in Fig.
 4(e) of the main text.
@@
 4(e) of the main text, the experimental tip height pm is aligned to the tip–sample distance nm for the calculations.
-Although the accurate determination of the experimental tip–sample gap distance is difficult, this alignment is reasonable; as described in Sect.
+Although the accurate determination of the experimental tip–sample gap distance is difficult, this alignment is reasonable; as described in Sec.
 I-E, pm is comparable with pm and nm.
 As shown in Fig.
-S11(a), the frequency and width of the H2 and D2 “rot” peaks are almost unchanged by the tip.
+S12(a), the frequency and width of the H2 and D2 “rot” peaks are almost unchanged by the tip.
 In contrast, the bottom panel Fig.
-S11(b) shows the strong peak broadening for the H2 “vib” mode [see also Fig.
-S10(b)] whereas the peak width for D2 “vib” is almost unchanged.
+S12(b) shows the strong peak broadening for the H2 “vib” mode [see also Fig.
+S11(b)] whereas the peak width for D2 “vib” is almost unchanged.
 A possible origin of the peak broadening is the anharmonicity, which can modify the lifetime of molecular vibrations [64].
@@
 2(c) in the main text], they would lead to the high-frequency shoulder of the vibration transition for H2, as these molecules would still lie on the surface [Fig.
-3(c)] and would therefore present a higher than the ones on the tip, which instead provide the most enhanced signal; however, this effect may not be pronounced because the coverage of is less than that of D2 (see Sect.
+3(c)] and would therefore present a higher than the ones on the tip, which instead provide the most enhanced signal; however, this effect may not be pronounced because the coverage of H2 is less than that of D2 (see Sec.
 I-C).